### v1

# Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese

Abstract
Large Language Models (LLMs) are increasingly being used to generate synthetic data for training and evaluating models. However, it is unclear whether they can generate a good quality of question answering (QA) dataset that incorporates knowledge and cultural nuance embedded in a language, especially for low-resource languages. In this study, we investigate the effectiveness of using LLMs in generating culturally relevant commonsense QA datasets for Indonesian and Sundanese languages. To do so, we create datasets for these languages using various methods involving both LLMs and human annotators. Our experiments show that the current best-performing LLM, GPT-4 Turbo, is capable of generating questions with adequate knowledge in Indonesian but not in Sundanese, highlighting the performance discrepancy between medium- and lower-resource languages. We also benchmark various LLMs on our generated datasets and find that they perform better on the LLM-generated datasets compared to those created by humans. [^1] All datasets and codes in this work are available in https://github.com/rifkiaputri/id-csqa. The dataset is available under the Creative Commons Non-Commercial (CC BY-NC 4.0).
Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese
[…]
1 Introduction
The development of Large Language Models (LLMs) is significantly impacting NLP, leading to an increasing trend in the automated generation of datasets, particularly for Question Answering (QA) tasks. However, a major challenge arises with non-English languages, especially those that are underrepresented, like Indonesian and Sundanese, due to the need for cultural context. For the generated data to be fully useful, it must not only be linguistically accurate, but it also needs to reflect the cultural nuances, historical references, and social norms. It is not yet clear whether current LLMs can create QA data that adequately includes the cultural nuances specific to languages such as Indonesian and Sundanese.
Figure 1: Our dataset generation methods. The examples of LLM_Adapt, Human_Gen, and LLM_Gen datasets are shown in English for clarity. The original versions of these datasets are in Indonesian and Sundanese.
In this study, we aim to investigate how well current LLMs generate a commonsense QA dataset that is both linguistically accurate and culturally relevant to Indonesia. We focus on Indonesian, the lingua franca of Indonesia, and Sundanese, one of the local languages in Indonesia. To sum up, our contributions are as follows:
  • We create a new Indonesian and Sundanese CommonsenseQA dataset using various methods, including adapting existing English data Talmor et al. (2019) and generating new datasets manually and automatically as illustrated in Figure 1. The dataset contains approximately 3K LLM-generated and 1.5K human-generated question-answer pairs per language, making it the largest of its kind. For human-generated data, we ensure cultural diversity by involving annotators from various regions, including Java and Bali.
  • We perform a comprehensive analysis to assess the effectiveness of LLMs in creating a culturally relevant commonsense QA dataset. We find that adaptation from English data is less effective, particularly for Sundanese. However, GPT-4 Turbo can generate questions with some basic local knowledge in Indonesian when provided with human-created categories and concepts. Yet, for Sundanese, questions often have major errors.
  • We evaluate several LLMs using our dataset and find that they perform better on the dataset generated by LLMs than humans. This finding reveals the limitation of the synthetic dataset that it is not as challenging for LLMs, especially for larger models such as GPT-4.
2 Related Work
Commonsense Dataset
Datasets and benchmarks for commonsense reasoning are crucial for developing models that understand and reason about real-world complexities. Datasets like COPA Roemmele et al. (2011), X-COPA Ponti et al. (2020), and The Winograd Schema Challenge Levesque et al. (2012) evaluate causal reasoning in real-world scenarios. Other datasets like ARC Clark et al. (2018), OpenBookQA Mihaylov et al. (2018), and Mcscript Ostermann et al. (2018) test commonsense reasoning via question-answering, but some questions require grade-school science knowledge. CommonsenseQA Talmor et al. (2019) presents the type of “purely” commonsense QA in the form of multiple-choice questions based on a concept knowledge from ConceptNet Speer et al. (2017), built through crowdsourcing. One question corresponds to one target concept, and other related concepts become distractors, acting as parts of the answer choices. However, since the dataset is crowdsourced, the crowdworkers bring their cultural background as part of their common knowledge, resulting in many questions that are “commonsense” only in the Western culture.
Cultural Evaluation Dataset
Various datasets are used to assess language models’ ability to understand cultural nuances specific to a language. FORK Palta and Rudinger (2023) explores culinary cultural biases and assumptions. However, this dataset is relatively small, comprising only 184 questions and focusing solely on US, Indian, and Chinese customs. In the context of Indonesian, IndoMMLU Koto et al. (2023) includes questions from Indonesian exams and covers regional cultural topics, such as the Minangkabau or Sundanese cultures. COPAL-ID Wibowo et al. (2023) is an Indonesian COPA-style dataset written by native speakers, thus incorporating more Indonesian cultural nuances compared to X-COPA. COPAL-ID is composed of approximately 500 questions, with a primary emphasis on the cultural aspects of the Jakarta region. In contrast, our dataset is much bigger and covers a broader range of annotators from various regions of Java and Bali, in addition to Jakarta. We also include Sundanese, addressing the gap in this low-resource language.
3 Background
3.1 Commonsense QA: Definition and Scope
The Cambridge Dictionary defines commonsense as "the basic level of practical knowledge and judgment that we all need to help us live in a reasonable and safe way." van Holthoon and Olson (1987) describe commonsense as "knowledge, judgment, and taste which is more or less universal and which is held more or less without reflection or argument." Both definitions highlight commonsense as the practical knowledge in daily life. Talmor et al. (2019) did not explicitly define the scope of commonsense in their QA dataset. Nevertheless, they outlined various skills required to answer the questions in their dataset correctly, such as spatial, cause & effect, social, activity, and many more. To build our dataset, we follow similar skills but focus on adding common or cultural knowledge specific to Indonesian and Sundanese contexts to reduce bias towards Western culture. Generally, one data point is a triple of concept, question, and options (with one correct answer). The data samples are shown in Figure 1.
3.2 Languages in Indonesia
Indonesia is one of the most culturally and linguistically diverse countries in the world, with more than 700 languages spoken across the country Aji et al. (2022); Eberhard et al. (2021). Of the many languages in Indonesia, Indonesian is a unifying language used nationally. It utilizes the Latin script and was developed from literary "Classical Malay" of the Riau-Johor sultanate Sneddon (2003), with regional variations. Over 80% of Standard Malay’s vocabulary is similar to Indonesian.
Apart from Indonesian, regional languages like Sundanese are spoken by people of the same ethnicity. Sundanese is primarily spoken in West Java, as well as Banten, Jakarta, Central Java, and more. [^2] https://petabahasa.kemdikbud.go.id/infobahasa2.php?idb=56&idp=Jawa%20Barat Regional languages including Sundanese have influenced the formation and development of the Indonesian language. Both languages share similarities, such as their grammatical structure, but also differ significantly in aspects like the number of vowels and morphological features, including affixes. We chose to study the national language and one regional language to illustrate the differences in the commonsense QA data generated via LLMs.
4 Data Generation Methods
To investigate whether LLMs can generate culturally relevant commonsense QA data in Indonesian and Sundanese, we construct a dataset using various methods with LLMs as data generators. In addition, we employ humans to generate data for comparison. As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation, where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation, where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation, where we utilize LLMs to generate data based on the list of categories and concepts used in method (2).
4.1 Automatic Data Adaptation
We build the first LLM-generated data by adapting the CommonsenseQA dataset Talmor et al. (2019) to make it culturally relevant to Indonesian and Sundanese. We employ a multi-stage approach to adapt the data. First, we remove data that contains ambiguous options and offensive keywords. Next step involves assessing three elements: concept, name, and location, in order to sample the data that will be adapted. We take the question concepts from the existing CommonsenseQA data. As for the name and location extraction, we use Stanford’s Stanza NLP library Qi et al. (2020) and ConceptNet Speer et al. (2017) API. We use GPT-3.5 Turbo results ensembled from five different prompts to determine the relevance of the data based on the three elements. Data that are considered irrelevant in at least one of these three elements are selected to be adapted.
The next step is to transform the selected irrelevant data. First, we prompt GPT-4 Turbo [^3] Our initial experiment (Appendix A.2) reveals that GPT-4 Turbo significantly outperforms Merak-v4, an open Indonesian LLM, leading us to select GPT-4 Turbo for our work. to rephrase the sampled questions and options to align with Indonesian cultures. Subsequently, for data flagged with Western-centric names, a second rephrasing is conducted using the GPT-3.5 Turbo to replace all person names. The fully rephrased and adapted data are then translated from Indonesian to Sundanese using Google Translation API. Finally, we filter the data by examining the appearance of concepts in the questions and checking Sundanese-Indonesian back-translation with a similarity threshold of 90%.
4.2 Manual Data Generation
We employ annotators with different ethnicities from diverse regions covering Java and Bali island to ensure the variability of the knowledge in the human-generated dataset. The data collection process involves two main phases: (1) creating commonsense question-answer pairs and (2) answering commonsense questions. We perform quality control with additional human annotators to ensure the data quality.
Creating Question-Answer Pairs
We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept. We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLM as much as possible. The categories and question concepts are manually curated to ensure wide domain coverage and to enrich the diversity of commonsense knowledge in the data. In total, we cover five categories: (1) culinary, (2) place, (3) culture, (4) history, and (5) activity, with 150 predefined question concepts for each category. Two annotators use the same 50 question concepts from each category to generate the questions, options, and answers. This process produces 1,500 QA pairs for each language in total.
Answering Question
After passing the quality assurance, the 1,500 QA pairs are then redistributed among all annotators for further review, where they are tasked with answering 1,250 commonsense questions each (excluding the set of data they have made in the first phase). This phase aims to evaluate the consistency and variance in the commonsense knowledge among annotators from various cultural backgrounds. Annotators are instructed not only to provide answers to the questions but also to comment on any ambiguities in the questions or options, or if they have any uncertainties when answering the questions.
Quality Control
We conduct Quality Control (QC) to ensure data correctness through manual human validation. QC annotators manually review data for errors and provide feedback, which is then corrected by the annotator concerned. Once corrected, QC annotators re-check the data to determine whether it can be considered complete or still requires revision. An evaluation meeting with all annotators is also held to convey a more comprehensive evaluation regarding all types of errors and other findings, ensuring annotators have a better understanding and do not repeat similar errors. After the second phase (answering questions), we also excluded some questions if more than three annotators marked them as ambiguous.
4.3 Automatic Data Generation
For generating the second type of LLM-generated data, rather than adapting questions from the English dataset (§4.1), we use the same set of categories and question concepts as the human-generated data (§4.2). We also utilize GPT-4 Turbo and instruct it to generate questions, options, and answers. To ensure the generated dataset aligns closely with the intended cultural context, we explicitly incorporated the categories and question concepts in the prompt. Additionally, we instruct the model to strictly include the question concepts in the generated questions. We adopt a batching approach to streamline the data generation process, providing the model with a maximum of 5 distinct question concepts from the same categories in one API call. To maintain dataset quality and uniqueness, we filter out duplicate entries and questions that do not explicitly contain the question concepts.
Dataset Version Indonesian Sundanese
Train Valid Test Train Valid Test
LLM_Adapt 1,506 191 158 1,506 191 158
Human_Gen - - 1,498 - - 1,499
LLM_Gen - - 1,063 - - 1,183
Total 1,506 191 2,719 1,506 191 2,840
Table 1: Statistics of our generated Indonesian and Sundanese CommonsenseQA dataset.
4.4 Resulting Data
Our final CommonsenseQA dataset contains 4,416 QA pairs for Indonesian and 4,537 QA pairs for Sundanese. Table 1 provides detailed statistics for both languages. We also have three data variations: (1) LLM_Adapt: LLM-generated dataset constructed through data adaptation method (§4.1); (2) Human_Gen: Human-generated dataset, as explained in §4.2; (3) LLM_Gen: LLM-generated dataset given the same set of categories and concepts in Human_Gen4.3).
5 Data Analysis
5.1 LLM-Generated Data
To evaluate the quality of our LLM-generated data, we manually reviewed all 158 samples from the LLM_Adapt test set. Based on the evaluation, we then calculate the accuracy of the generated concepts, questions, and options. Furthermore, we evaluate 300 randomly selected samples from the LLM_Gen using the same procedure.
5.1.1 Concept Analysis
Concept Quality
In the English (eng) to Indonesian (ind) concept adaptation, 95.57% are correctly adapted, including 16 out of 19 concepts needing major adaptation for cultural relevance, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall). Most adaptations are minor, primarily involving direct translations of eng to ind, as shown in Table 2. For Indonesian to Sundanese, the correct adaptation rate drops to 77.22%, reflecting weaker machine translation performance for Sundanese. The samples are shown in Table 3.
Status Num (%) Concept Example
Orig (eng) Modified (ind)
Correct (major) 32 (20.25%) beaver
komodo
snow
hujan abu vulkanik
(volcanic ashfall)
Correct (minor) 119 (75.32%) tower menara (tower)
grape anggur (grape)
Wrong 7 (4.43%) orchestra pit
sumur orkestra
(orchestra well)
skate ice skating
Table 2: Samples of English (eng) to Indonesian (ind) concept adaptation.
Status Num (%) Concept Example
Orig (ind) Modified (sun)
Correct 122 (77.22%)
hujan abu vulkanik
(volcanic ashfall)
hujan lebu vulkanik
(volcanic ashfall)
menara (tower) munara (tower)
Wrong 36 (22.78%)
cicak
(house gecko)
kadal imah
(house lizard)
klinik gigi
(dental clinic)
klinik dental
(dental clinic)
Table 3: Samples of Indonesian (ind) to Sundanese (sun) concept adaptation.
Concept Variation
Despite the high accuracy of concept adaptation, as shown in Figure 2, we observe that the adapted concepts are skewed towards ‘komodo’. This indicates that the model is biased toward a specific entity within a category, which could arise from its limited knowledge or the absence of a direct equivalent for the original English concept in Indonesian, leading it to default to one standard concept. This finding highlights the need to manually develop concepts across various categories to improve the dataset’s quality in terms of concept diversity and coverage.
Figure 2: Top-10 adapted question concepts taken from train, validation, and test set of LLM_Adapt data.
5.1.2 Question Analysis
Question Quality
To evaluate the quality of the generated questions, we apply a strict criterion: any errors, even minor, are marked as incorrect. As shown in Table 4, the Indonesian datasets show a high percentage of error-free questions, ranging between 68–75%. However, for Sundanese, the accuracy decreases significantly. In particular, the weak performance of the MT system in Sundanese is evident from its very low accuracy in LLM_Adapt. This suggests that synthetic data adaptation and generation is not useful for low-resource languages.
Common Mistakes
To further analyze the quality of the questions, we delve deeper into the mistakes that are commonly made, which are listed in Table 5. We observe that most of the errors in LLM_Adapt come from translation errors. For LLM_Gen, even though the accuracy of question generation is lower than LLM_Adapt, 54.26% of the errors in LLM_Gen are actually minor typos or mechanics errors. In the case of Sundanese, most errors in LLM_Gen are due to sentence fluency, indicating that the model is still struggling to generate fluent sentences in Sundanese.
5.1.3 Options and Answer Analysis
Options Quality
We also evaluate the options in a similar manner to the evaluation of the question. From Table 6, we observe that LLM_Gen generates higher quality options compared to LLM_Adapt. However, there is still a significant gap in performance between Indonesian and Sundanese. This once again highlights the performance discrepancy between medium- and lower-resource languages.
Common Mistakes
We conduct a detailed analysis to identify the common errors found in each dataset, shown in Table 7. For LLM_Adapt, the most common issues are typos and mechanical errors, particularly with capitalization (e.g., "indonesia" instead of "Indonesia"), along with translation errors. In the Sundanese dataset, the errors are mainly due to major translation issues. As for LLM_Gen, although errors constitute less than 10% of the total evaluated samples, the primary issue is the presence of invalid options, particularly when no correct answers are among the options. For Sundanese, alongside translation errors, the model also produces some invalid options.
Dataset % of correct questions
ind sun
LLM_Adapt 75.32% 15.19%
LLM_Gen 68.67% 51.00%
Table 4: Question generation accuracy of LLM-generated datasets. The accuracy reflects the percentage of a ‘clean’ question without any major or minor errors.
Error Type % of error type
LLM_Adapt LLM_Gen
ind sun ind sun
Translation 33.33% 48.51% 0.00% 13.61%
Sent. structure 12.82% 11.94% 0.00% 0.00%
Sent. fluency 28.21% 27.61% 36.17% 36.73%
Sent. context 7.69% 2.24% 9.57% 16.33%
Subjectivity 2.56% 0.75% 0.00% 0.00%
Typo/mechanics 15.38% 8.96% 54.26% 33.33%
Table 5: Distribution of question generation error types of LLM-generated datasets.
Dataset % of correct choices
ind sun
LLM_Adapt 62.66% 38.61%
LLM_Gen 93.00% 58.67%
Table 6: Choices generation accuracy of LLM-generated datasets. The accuracy reflects the percentage of ‘clean’ choices.
Error Type % of error type
LLM_Adapt LLM_Gen
ind sun ind sun
Translation 10.17% 77.32% 0.00% 76.61%
Sent. fluency 1.69% 1.03% 9.52% 3.23%
Sent. context 1.69% 1.03% 0.00% 0.00%
Invalid options 3.39% 2.06% 80.95% 20.16%
Typo/mechanics 83.05% 18.56% 9.52% 0.00%
Table 7: Distribution of choices generation error types of LLM-generated datasets.
5.2 LLM vs. Human-Generated Data
To compare the data generated by LLM with human-generated data, we analyze the proportion of shared unigram and bigram tokens between the LLM_Gen and Human_Gen datasets. The overlap percentage is calculated by dividing the number of shared tokens by the total number of tokens. We find that the unigram overlap percentage is 39.75% for Human_Gen and 65.48% for LLM_Gen. A similar trend is observed for the bigram overlap percentage, with 12.41% for Human_Gen and 15.98% for LLM_Gen. This shows that many tokens present in Human_Gen also exist in LLM_Gen, but the reverse is not equally true. Given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol. [^4] In the context of culinary, ‘kalis’ means a state of dough that is well-kneaded and ready to be processed further. ‘Cimol’ is an Indonesian street food made from tapioca flour.
Upon sample-level analysis, we also find token variance disparities, particularly related to cultural context. For instance, given ‘kerupuk’ (crackers) concept, human annotators can formulate questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers). In contrast, LLM tends to ask more general questions, like "What is the common primary ingredient of crackers?" Also, although LLM can generate some culturally related entities such as ‘tari kecak’, a traditional dance from Bali, it is more popular compared to ‘tari pendet’, which is generated by the human annotator. More examples are outlined in Table 11.
6 Benchmark Result
6.1 Experiment Setup
We conduct a zero-shot evaluation of various LLMs to assess their performance on our datasets.
English-centric LLMs
We include LLaMA-2 7B and 13B Touvron et al. (2023), a widely used open LLM, and MistralOrca-7B Lian et al. (2023).
Multilingual LLMs
We include PolyLM-13B Wei et al. (2023), an LLM trained on a multilingual dataset, predominantly in English and Chinese; BLOOMZ-7B Muennighoff et al. (2022), an LLM fine-tuned on the xP3 dataset; and SeaLLM-7B Nguyen et al. (2023), an LLM with a focus on Southeast Asian languages.
Monolingual LLMs
We include Merak-v4 Ichsan (2023), an Indonesian LLM fine-tuned based on MistralOrca-7B. We also include MalaysianMistral-7B Zolkepli et al. (2024), a Malaysian LLM extended from Mistral-7B.
Proprietary LLMs
We include GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo, OpenAI (2023), which are accessed via OpenAI API.
We use three prompt variations for all models (see Appendix E). For a fair comparison between open and proprietary LLMs, we extract the answer key from the text generation result instead of the next token probability, using a rule-based and regex. Evaluations are performed on RTX A6000 48GB. We use accuracy as the evaluation metric.
6.2 Overall Performance
Figure 3: Overall performance of LLMs on our combined test set. We run three different prompts for each question in a zero-shot setting.
[…]
6.3 LLM vs. Human-Generated Data
Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen.
To assess LLMs’ ability to answer both LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets. Our result, shown in Figure 4, reveals that LLMs perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data. This indicates that the model may possess some knowledge to answer LLM-generated questions, even though it was not specifically trained on Indonesian or Sundanese data. However, the model still struggles to answer new knowledge present in Human_Gen data. The performance gap is also observed in Indonesian LLM, Merak-v4. The accuracy of this model in Sundanese is also low, indicating that training LLM solely on Indonesian texts is not enough to make it perform well in local languages such as Sundanese.
6.4 Performance by Question Category
Figure 5: LLMs performance by question category in LLM_Gen and Human_Gen for Indonesian and Sundanese.
Figure 5 shows the LLMs’ performance across different question categories. It reveals that LLMs perform better in the activity and place questions, but struggle in culinary. A notable performance gap exists between LLM_Gen and Human_Gen dataset, with weaker results on the latter. Moreover, open LLMs exhibit a significant performance gap compared to proprietary models, especially in Human_Gen, indicating their shortcomings in capturing commonsense knowledge. Interestingly, a significant improvement is seen from GPT-4 (gpt-4-0613) to GPT-4 Turbo (gpt-4-1106), especially in Sundanese Human_Gen. This suggests that the model is acquiring more knowledge, one reason for which possibly comes from the interactions on the ChatGPT web. Despite these advancements, some categories still score below 80%, showing there is still room for improvement.
7 Discussion
7.1 Multiple-Choice vs. ‘Free’ Generation
Since our dataset is in a multiple-choice format, LLMs might look better than they actually are because they can just pick one answer from the given options. To see LLMs’ genuine capability, we test GPT-4 Turbo on 100 randomly sampled questions from Indonesian Human_Gen in open-ended settings, i.e., asking the question directly without showing answer options or providing any extra instructions. Our manual evaluation shows a 77% accuracy rate, highlighting the model’s struggle with open-ended questions and confirming our dataset’s importance. Among the wrong answers, 18 cases stemmed from overly general answers not specific to Indonesia. For example, when asked "Where can we find many villas?" the model answers "Bali" but also listed other cities in mostly Western countries. In contrast, the annotator writes "Bandung" as the correct answer. As for the other wrong cases, the model struggles the most in answering questions in the culinary category. More failure examples can be seen in Table 12.
7.2 Extension to Other Local Languages
Besides Indonesian, our study focuses on one Indonesian local language, Sundanese. As previously discussed, Indonesia has many local languages; however, we cannot cover all of them due to resource constraints. We aim for our findings in Sundanese to act as a starting point for other languages. We anticipate that LLMs might perform worse than Sundanese for extremely low-resource languages, like Buginese or Toba Batak. Javanese, on the other hand, is expected to have comparable performance to Sundanese Winata et al. (2023); Bang et al. (2023). Generating good-quality synthetic data could be particularly beneficial for extremely low-resource languages, such as Buginese or Toba Batak, given the difficulty of finding native speakers. Yet, our case study in Sundanese indicates that LLMs might struggle even more with understanding cultural nuances in such languages.
8 Conclusion
In this study, we created Indonesian and Sundanese CommonsenseQA datasets using various dataset collection methods, including automatic dataset generation utilizing LLMs and manual dataset generation employing human annotators from diverse regions. Our thorough analysis reveals a substantial gap in the performance of LLMs between Indonesian and Sundanese, both in the dataset generation and in answering the questions in the dataset itself. Thus, one must be cautious when using existing LLMs in generating datasets, particularly for lower-resource languages, since their culturally relevant commonsense knowledge may be limited. Our findings highlight the urgent need to create high-quality resources for these underrepresented languages to prevent them from being left behind.
Limitations
Our work has some limitations that should be acknowledged. Firstly, in terms of language coverage, we were only able to cover Indonesian and Sundanese due to the available resources and the authors’ familiarity with these languages. Additionally, the annotators we recruited were mostly from Java island, with one annotator from Bali island. Despite our effort to include a range of question concepts from different regions, including those beyond Java and Bali islands, it is possible that some bias may exist, especially in the Indonesian dataset. This is because the questions were generated primarily by annotators from Java and Bali, and their perspectives and cultural backgrounds may have influenced the content. Nonetheless, we have taken measures to eliminate potentially harmful questions.
Additionally, we use a multiple choice question format, following English CommonsenseQA data format Talmor et al. (2019) to facilitate a more straightforward and robust evaluation process. Although open-ended generation may offer a more challenging benchmark for LLMs, as discussed in Section 7.1, evaluating LLMs in such settings poses its own set of challenges, especially in low-resource languages where ‘LLM-as-a-judge’ approach may not be as effective as in English. Nevertheless, our dataset can serve as a starting point for this line of research direction.
Ethical Consideration
All human-generated datasets have been manually validated to ensure that harmful or offensive questions are not present in the dataset. We also excluded potentially harmful questions in the LLM-generated datasets through automatic filtering. Our work has been reviewed by the Institutional Review Board (IRB) in our institution. All recruited annotators were paid above the minimum wage. Our datasets will be publicly available under the Creative Commons Non-Commercial (CC BY-NC 4.0) license.
Appendix A LLM-Generated Data
A.1 Prompts Details
A.1.1 Automatic Data Adaptation
There are two prompts used to adapt the CommonsenseQA into LLM_Adapt: Adapt All and Adapt Name. Below are the details of both prompts.
Adapt All Prompt
[…]
[leftmargin=0pt, rightmargin=0pt]
Change the given data to make it relevant to Indonesia in any ways. Make all elements relevant to each other, and the concept always appear explicitly in the question. Return in Indonesian language with JSON format where question is string, concept is string, options is dictionary where label is the keys and option text is the values, and question_answer is string contain one label from the options.
Data:
Question: {QUESTION}
Concept: {QUESTION CONCEPT}
Options:
{CHOICES}
Question Answer: {ANSWER}
[…]
Adapt Name Prompt
[…]
[leftmargin=0pt, rightmargin=0pt]
Change all names in the given question to Indonesian names. Change only the names, keep all other phrases in the question the same and keep it all in Indonesian.
Question: {QUESTION}
Changed Question:
[…]
A.1.2 Automatic Data Generation
For LLM_Gen, we directly generate the data given a set of question concepts. Below are the prompt used for data generation.
Data Generation Prompt
[…]
[leftmargin=0pt, rightmargin=0pt]
Given a list of {LANGUAGE} concepts [QUESTION CONCEPTS}], create one {LANGUAGE} commonsense QA data with topic "{CATEGORY}" for each concept, that consists of three components: "question", "choices", and "answer_creator". The "question" must contains the concept explicitly. The "choices" consist of 5 different choices marked A to E where one should be the "answer_creator". All data should be in {LANGUAGE}, return only your answer in JSON data format, and add the concept of the data as "question_concepts".
JSON Data:
[…]
A.2 Choosing Model for Data Generation: Indonesian LLM vs. GPT-4 Turbo
Model Win Rate
Concept Question Choices
Automatic Data Adaptation
Merak-v4 28.0% 8.5% 4.0%
GPT-4 Turbo 72.0% 91.5% 96.0%
Automatic Data Generation
Merak-v4 - 10.5% 9.5%
GPT-4 Turbo - 89.5% 90.5%
Table 8: Win rate comparison of Merak-v4 (open Indonesian LLM) and GPT-4 Turbo (best-performing proprietary LLM).
Table 8 shows the win rate accuracy of concept, question, and choices generation of Indonesian LLM, Merak-v4, and best-performing proprietary LLM, GPT-4 Turbo. The results demonstrate that GPT-4 Turbo significantly outperforms Merak-v4, with win rates ranging from 72% to 96%. Our sample-level analysis indicates that a significant number of questions generated by Merak-v4 tend to have obvious answers or are formulated as yes/no questions, such as "Apakah jagung dapat dimakan?" (eng: Is corn edible?). Additionally, despite being trained on Indonesian texts, the model occasionally produces questions with US-centric knowledge. For instance, it generates "Siapa yang memimpin kampanye pemilihan presiden pertama di Amerika Serikat?" (eng: Who led the first presidential election campaign in the United States?).
A.3 Additional Analysis of the Generated Data: Common Mistakes in Concept Adaptation
Error Type Num (%) of errors
eng normal-$\rightarrow$ ind ind normal-$\rightarrow$ sun
Translation 4 (57.14%) 31 (86.11%)
Phrase structure 1 (14.29%) 1 (2.78%)
Typo/mechanics 2 (28.57%) 4 (11.11%)
Table 9: Summary of errors in concept adaptation.
The details of common mistakes in concept adaptation for Indonesian and Sundanese are shown in Table 9. In the case of adapting concepts from eng to ind, many of the errors are translation errors, resulting from awkward phrasing of the translated concept or the concept remaining in English instead of being translated to Indonesian. However, the number of errors is relatively small. In the case of adapting concepts from ind to sun, similar to eng to ind, the majority of errors also arise from translation errors, with a larger number of errors.
Appendix B Human Annotation Guideline
To ensure a high-quality and standardized format for question-answers annotation, we provide a specific guideline during the annotation process. The process of creating question-answer pair data is carried out using Google Spreadsheets. Each person will get a Google Sheets document that will be their worksheet. The explanation of each field is described below.
  1. 1.
    ID: This column contains the ID of each data.
  2. 2.
    Category: This column contains the overarching category of the data.
  3. 3.
    Question Concept: This column contains concepts from categories that need to be used in creating questions.
  4. 4.
    Question: This column is used to write commonsense questions that contain the given concept and match the category.
  5. 5.
    Choices: This column is used to write 5 choices for the questions given. Of the 5 choices given, the annotator needs to ensure there is 1 correct answer and 1 distractor. Distractor is an answer that could potentially be considered the correct answer.
  6. 6.
    Answer: This column only needs to be filled in when the row contains the correct answer.
  7. 7.
    Distractor: This column only needs to be filled in when the row contains the distractor.
Categories and Concepts
Five selected categories are used in the data creation process. Each category has 150 concepts that need to be used and developed to become commonsense questions. The following is a list of the categories.
  1. 1.
    Culinary: Concepts in this category include everything related to culinary, starting from culinary types, cooking ingredients, cooking tools & methods, to etiquette or eating habits.
  2. 2.
    Places: Concepts in this category include everything related to places, starting from public facilities, landmarks, buildings, and various other concepts related to places.
  3. 3.
    Culture: Concepts in this category include everything related to culture, starting from cultural elements, cultural tools, cultural actors, to customs and habits that exist in Indonesia.
  4. 4.
    History: Concepts in this category include everything related to history, starting from historical events, historical actors, historical findings, and various other concepts related to history.
  5. 5.
    Activities: Concepts in this category include everything related to activities, starting from sports, hobbies, household work, and various other concepts related to activities.
General Rules
There are several rules in the data creation process. The following is the explanation.
  1. 1.
    The questions asked are commonsense questions (not factual questions) related to Indonesian/Sundanese culture. Especially for the History category, questions can also be in the form of factual questions, but they must be general facts (commonly known).
  2. 2.
    The questions created must be related to the given category and must contain the given concept.
  3. 3.
    Annotators are allowed to change the morphological form of concepts as long as they do not change the categories and basic words. (i.e. berkunjung $\rightarrow$ kunjungan, mengunjungi, dikunjungi).
  4. 4.
    The priority is that the questions asked are always related to general daily life or Indonesian/Sundanese culture.
  5. 5.
    Both language groups will use the same list of categories and concepts. However, when creating data, annotators are expected to incorporate perspectives that align with their respective cultural backgrounds.
  6. 6.
    Each person will get 50 concepts from each category. The total number of concepts that will be accepted is 250 concepts, equivalent to the workload for each person, where the expectation is to create one question per concept.
Appendix C Annotators Demographics
Data Team Ethnicity Domicile Num
Indonesian Sundanese Sukabumi, West Java 1
Depok, West Java 1
Javanese Magelang, Central Java 1
Bojonegoro, East Java 1
Betawi Tangerang, Banten 1
Balinese Denpasar, Bali 1
Sundanese Sundanese Bandung, West Java 2
Bogor, West Java 1
Majalengka, West Java 1
Sukabumi, West Java 2
Total 12
Table 10: Demographic information of the annotators from each dataset team. Note that even though Depok is included in West Java Province and Tangerang is included in Banten Province, both are geographically closer to Jakarta and considered as part of the Greater Jakarta area (Jabodetabek).
In accordance with the aim of constructing the dataset, we involved all Indonesian native-speaker annotators in the manual data generation process. All the annotators who worked on the Indonesian dataset were people from several regions on the islands of Java and Bali. Meanwhile, the annotators involved in building the Sundanese dataset were Sundanese people who come from several different regions in West Java. Each dataset development team consists of 6 annotators. The detailed annotators’ demographics are shown in Table 10.
Appendix D Manual Data Generation Findings
During the manual data generation process, we discovered several interesting findings from the data created by the annotators.
Questions
We found quite a lot of interesting findings in the process of creating commonsense questions. We found that there were questions created using very general contexts, while the limitations of this data generation were the Indonesian and Sundanese contexts (specifically for the Sundanese dataset). For instance, "Apa jenis restoran yang paling terkenal di seluruh dunia?" (eng: Which type of restaurant is most famous worldwide?) In that question, the annotator developed the concept question “restoran” (eng: restaurant) into a question with a very general context: “di seluruh dunia" (eng: all over the world). We also found subjective questions, so the answers that emerged also had subjective value (not commonsense). These subjective questions are usually characterized by the use of superlative adjectives. Another type of finding related to the question category is the creation of logical/causal questions. This seems to be based on confusion from annotators regarding the boundaries of commonsense and logic/causation.
Choices
The findings related to choices that are most often encountered are making choices that have the same value as each other so that it is difficult to determine the answer and distractor. For instance, the question "Kue apa yang biasanya disajikan pada momen lebaran?" (eng: What cakes are usually served during Eid?). All the options given are types of cakes that are generally served during Eid. This finding is also related to the issue of subjectivity.
Answers-Distractors
Answers and distractors are another category that also has many findings in the manual data generation process. This issue is not much different from the findings in Choices: answers and distractors are equivalent. Apart from that, in this category, it was also found that personal experience was used in determining answers so that more common answers were determined as distractors. For instance, the question "Apa yang biasanya orang lakukan di stasiun?" (eng: What do people usually do at the train station?). The annotator specified "mengantar teman/kerabat" (eng: accompany friends/relatives) as the answer. While the "naik-turun kereta" (eng: get on and off the train) option was chosen as a distractor.
Knowledge Variations between Annotators
Figure 6: Answer conflict across Indonesian annotators.
Figure 7: Answer conflict across Sundanese annotators.
As we employ annotators from different regions, it enables us to explore variations in the data they generate. We examine this by calculating the number of answer conflicts that arise during the "answering question" phase of our data generation pipeline (§4.2). The results from Indonesian and Sundanese annotators are detailed in Figure 6 and 7, respectively. Our analysis revealed that, out of the Indonesian annotators, the one from Bali has the highest number of conflicting answers. However, the number of questions generated by the Bali annotator does not seem to have a lot of conflicts, suggesting that the annotator tends to generate easier questions, some of which have an obvious answer. This finding contrasts with the results from West Java annotators, who, despite generating questions that lead to a higher number of conflicts, do so mainly due to the creation of more challenging option distractors.
In the case of Sundanese annotators, the variation in answer conflicts across regions is not significant. This lack of variation can likely be attributed to the Sundanese language’s narrower geographic distribution which primarily spoken in West Java, unlike the Indonesian language, which serves as a lingua franca within the country. This finding points to the fact that differences in annotators’ knowledge levels may arise from various factors, one of which could be the regional culture to which they are most exposed.
Appendix E Zero-Shot Benchmark Prompts
We apply three prompt variations to test LLMs’ performance on our CommonsenseQA datasets. Each prompt is described below.
Prompt Variation 1
[…]
[leftmargin=0pt, rightmargin=0pt]
The following are multiple choice questions (with answers) about "{CONCEPT}".
{QUESTION}
A. {CHOICE_A}
B. {CHOICE_B}
C. {CHOICE_C}
D. {CHOICE_D}
E. {CHOICE_E}
Answer:
[…]
Prompt Variation 2
[…]
[leftmargin=0pt, rightmargin=0pt]
Question: {QUESTION}
Choices:
A. {CHOICE_A}
B. {CHOICE_B}
C. {CHOICE_C}
D. {CHOICE_D}
E. {CHOICE_E}
Answer:
[…]
Prompt Variation 3
[…]
[leftmargin=0pt, rightmargin=0pt]
The following are multiple choice questions (with answers) about "{CONCEPT}".
Question: {QUESTION}
A. {CHOICE_A}
B. {CHOICE_B}
C. {CHOICE_C}
D. {CHOICE_D}
E. {CHOICE_E}
Answer:
[…]
Figure 8: Performance comparison of raw vs. cleaned version of LLM-generated data.
Appendix F Effect of Synthetic Dataset Cleaning
To check how much cleaning the noise in synthetic (LLM-generated) data affects LLMs performance in answering our questions, we manually correct the errors in all 158 test sets of LLM_Adapt and 300 randomly sampled data from the LLM_Gen. The results are presented in Figure 8. The performance variance appears smaller on the cleaned dataset for LLM_Adapt, leading to more consistent performance. However, the performance improvement is not as significant for the LLM_Gen dataset, likely due to the lesser noise in this dataset compared to LLM_Adapt. Thus, a robust model capable of producing cleaner data can minimize the need for extensive dataset cleaning. However, caution is still necessary as errors can still occur.
Category Concept Data Creator Question Choices Answer
activity panen (harvest) GPT-4 Turbo
Kegiatan apa yang biasanya dilakukan
ketika musim panen tiba?
(What activity is usually done when
the harvest season arrives?)
A. Makan bersama keluarga
(eating together with family)
B. Berlibur (going on a holiday)
C. Memetik hasil pertanian
(picking agricultural produce)
D. Memancing (fishing)
E. Berbelanja (shopping)
C
Human
Kegiatan apa yang masyarakat desa lakukan
setelah panen mereka berhasil?
(What activity do village communities do
after their harvest succeeds?)
A. Bercocok tanam (farming)
B. Memotong rumput
(cutting grass)
C. Syukuran
(having a ‘syukuran’ event)
D. Jalan-jalan bersama
(going out together)
E. Wortel (carrots)
C
culinary kerupuk (crackers) GPT-4 Turbo
Kerupuk biasanya terbuat dari
bahan dasar apa?
(What is the common primary
ingredient of crackers?)
A. Tepung terigu (wheat flour)
B. Tepung beras (rice flour)
C. Tepung tapioka (tapioca flour)
D. Tepung jagung (corn flour)
E. Tepung kentang (potato flour)
C
Human
Terbuat dari apakah kerupuk rambak?
(What is ‘rambak’ crackers made of?)
A. Kulit hewan (animal skin)
B. Lidah hewan (animal tounge)
C. Usus hewan (animal intestines)
D. Nasi (rice)
E. Tepung beras (rice flour)
A
culture tarian (dance) GPT-4 Turbo
Tarian apa yang merupakan
tarian tradisional dari Bali?
(Which dance is a traditional
dance from Bali?)
A. Tari Pendet (Pendet dance)
B. Tari Saman (Saman dance)
C. Tari Kecak (Kecak dance)
D. Tari Jaipong (Jaipong dance)
E. Tari Yapong (Yapong dance)
C
Human
Manakah yang termasuk tarian dari Bali?
(Which includes a dance from Bali?)
A. Tari Piring (Piring Dance)
B. Tari Pendet (Pendet Dance)
C. Tari Topeng (Topeng Dance)
D. Tari Sajojo (Sajojo Dance)
E. Tari Melasti (Melasti Dance)
B
history era (era) GPT-4 Turbo
Era apa yang ditandai dengan berakhirnya
Perang Dingin dan runtuhnya Uni Soviet?
(Which era is marked by the end of the
Cold War and the collapse of
the Soviet Union?)
A. Era Globalisasi
(globalization era)
B. Era Informasi (information era)
C. Era Reformasi (reformation era)
D. Era Pasca-Perang Dingin
(pasca-cold war era)
E. Era Industrialisasi
(industrialization era)
D
Human
Berapa lama era Orde Baru
berlangsung di Indonesia?
(How long did the New Order era
last in Indonesia?)
A. 4 tahun (4 years)
B. 5 tahun (5 years)
C. 20 tahun (20 years)
D. 25 tahun (25 years)
E. 32 tahun (32 years)
E
place gunung (mountain) GPT-4 Turbo
Di manakah biasanya gunung berada?
(Where are mountains usually located?)
A. Di kota (in the city)
B. Di dataran tinggi (in highlands)
C. Di dalam laut (in the sea)
D. Di padang pasir (in the desert)
E. Di pinggir sungai
(at the riverbank)
B
Human
Apa nama gunung tertinggi
di Indonesia?
(What is the name of the
highest mountain in Indonesia?)
A. Gunung Jaya Wijaya
(Mt. Jaya Wijaya)
B. Gunung Bromo (Mt. Bromo)
C. Gunung Merapi (Mt. Merapi)
D. Gunung Semeru (Mt. Semeru)
E. Gunung Sumbing (Mt. Sumbing)
A
Table 11: Examples of data generated by GPT-4 Turbo (gpt-4-1106-preview) compared to human annotators.
Appendix G LLM vs. Human-Generated Data Examples
We provide some GPT-4 Turbo generation samples compared with human-generated data in Table 11.
Prompt Response Failure Reason
Apa judul lagu yang wajib dinyanyikan pada sesi mengheningkan cipta di setiap upacara bendera?
(What is the mandatory song to be sung during the ‘mengheningkan cipta’ (moment of silence) in every flag ceremony?)
Pada sesi mengheningkan cipta di setiap upacara bendera di Indonesia, biasanya tidak dinyanyikan lagu. …
(During the moment of silence in every flag ceremony in Indonesia, a song is usually not sung. …)
The mandatory song to be sung during the ‘mengheningkan cipta’ (moment of silence) session in flag ceremonies in Indonesia is ‘Hymne Pahlawan (Mengheningkan Cipta).’
Manakah menu yang dikonsumsi menggunakan sumpit?
(Which dishes are consumed with chopsticks?)
Sumpit adalah alat makan yang biasanya digunakan dalam tradisi kuliner Asia Timur, seperti Jepang, Korea, dan China. Berikut adalah beberapa contoh menu yang sering dikonsumsi menggunakan sumpit: Sushi dan Sashimi (Jepang), Ramen (Jepang), … [proceed with other dish examples in Japan, Korea, and China]
(Chopsticks are utensils that are commonly used in East Asian culinary traditions, such as Japan, Korea, and China. Here are some examples of dishes often consumed with chopsticks: Sushi and Sashimi (Japan), Ramen (Japan), … [proceed with other dish examples in Japan, Korea, and China])
Many Indonesian foods are commonly eaten with chopsticks, such as mie ayam (chicken noodles), bakmi, mie yamin, and many more.
Di daerah mana yang terdapat banyak vila?
(In which area can we find many villas?)
… beberapa daerah yang terkenal dengan banyaknya vila adalah: 1. Bali, Indonesia 2. Tuscany, Italia 3. Provence, Prancis [proceed with other location in Spain, US, Thailand, Portugal, US]
(… some areas known for having many villas include: 1. Bali, Indonesia 2. Tuscany, Italy 3. Provence, France [proceed with other locations in Spain, US, Thailand, Portugal, US])
Apart from Bali, there are many other areas in Indonesia famous for having many villas, such as Bandung, Lembang, Puncak, Batu, and many more.
Table 12: Examples of incorrect responses by GPT-4 Turbo (gpt-4-1106-preview) in ‘free’ or open-ended generation settings.
Appendix H Examples from Multiple-Choice vs. ‘Free’ Generation Experiment
We provide some GPT-4 Turbo generation samples when given the question in ‘free’ or open-ended generation settings in Table 12.

### v2

@@
 Abstract
-Large Language Models (LLMs) are increasingly being used to generate synthetic data for training and evaluating models.
-However, it is unclear whether they can generate a good quality of question answering (QA) dataset that incorporates knowledge and cultural nuance embedded in a language, especially for low-resource languages.
-In this study, we investigate the effectiveness of using LLMs in generating culturally relevant commonsense QA datasets for Indonesian and Sundanese languages.
-To do so, we create datasets for these languages using various methods involving both LLMs and human annotators.
-Our experiments show that the current best-performing LLM, GPT-4 Turbo, is capable of generating questions with adequate knowledge in Indonesian but not in Sundanese, highlighting the performance discrepancy between medium- and lower-resource languages.
-We also benchmark various LLMs on our generated datasets and find that they perform better on the LLM-generated datasets compared to those created by humans.
-[^1] All datasets and codes in this work are available in https://github.com/rifkiaputri/id-csqa.
-The dataset is available under the Creative Commons Non-Commercial (CC BY-NC 4.0).
-Can LLM Generate Culturally Relevant Commonsense QA Data? Case Study in Indonesian and Sundanese
 […]
+Can LLM Generate Culturally Relevant Commonsense QA Data?
+Case Study in Indonesian and Sundanese
+Rifki Afina Putri1, Faiz Ghifari Haznitrama1, Dea Adhista2, Alice Oh1 1KAIST, Republic of Korea 2Independent Researcher {rifkiaputri,haznitrama}@kaist.ac.kr, deadhista@gmail.com, alice.oh@kaist.edu
 1 Introduction
 The development of Large Language Models (LLMs) is significantly impacting NLP, leading to an increasing trend in the automated generation of datasets, particularly for question answering (QA) tasks.
-However, a major challenge arises with non-English languages, especially those that are underrepresented, like Indonesian and Sundanese, due to the need for cultural context.
+However, a major challenge arises with underrepresented languages like Indonesian and Sundanese due to the need for cultural context.
 For the generated data to be fully useful, it must not only be linguistically accurate, but it also needs to reflect the cultural nuances, historical references, and social norms.
-It is not yet clear whether current LLMs can create QA data that adequately includes the cultural nuances specific to languages such as Indonesian and Sundanese.
+It is not yet clear whether current LLMs can create QA data that adequately includes the cultural nuances specific to certain languages.
+Another common method for constructing non-English datasets is by utilizing a machine translation model.
+Although more scalable, this method cannot be straightforwardly applied due to the contextual irrelevancy of the data, primarily due to geographical differences (which could also influence cultural differences).
+For example, many English CommonsenseQA questions Talmor et al.
+(2019) include concepts such as snow or any winter sports, which are irrelevant in Indonesia due to its tropical climate with only two seasons.
+This dataset also often includes English-centric names and locations, primarily limited to the US, which are not considered “commonsense” for Indonesians.
 Figure 1: Our dataset generation methods.
@@
 The original versions of these datasets are in Indonesian and Sundanese.
-In this study, we aim to investigate how well current LLMs generate a commonsense QA dataset that is both linguistically accurate and culturally relevant to Indonesia.
-We focus on Indonesian, the lingua franca of Indonesia, and Sundanese, one of the local languages in Indonesia.
+Therefore, in this study, we aim to investigate how well current LLMs can adapt and generate a commonsense QA dataset that is both linguistically accurate and culturally relevant to Indonesia.
+We focus on Indonesian, the lingua franca of Indonesia, and Sundanese, one of the regional languages in Indonesia with 34 million speakers, yet still considered low-resource Aji et al.
+(2022).
+To address dataset scarcity, especially in Sundanese, we also manually constructed datasets by involving annotators from various regions in Indonesia, ensuring a representation of diverse cultural perspectives.
 To sum up, our contributions are as follows:
-  • We create a new Indonesian and Sundanese CommonsenseQA dataset using various methods, including adapting existing English data Talmor et al.
-(2019) and generating new datasets manually and automatically as illustrated in Figure 1.
-The dataset contains approximately 3K LLM-generated and 1.5K human-generated question-answer pairs per language, making it the largest of its kind.
-For human-generated data, we ensure cultural diversity by involving annotators from various regions, including Java and Bali.
+[…]
   • We perform a comprehensive analysis to assess the effectiveness of LLMs in creating a culturally relevant commonsense QA dataset.
-We find that adaptation from English data is less effective, particularly for Sundanese.
-However, GPT-4 Turbo can generate questions with some basic local knowledge in Indonesian when provided with human-created categories and concepts.
-Yet, for Sundanese, questions often have major errors.
-  • We evaluate several LLMs using our dataset and find that they perform better on the dataset generated by LLMs than humans.
-This finding reveals the limitation of the synthetic dataset that it is not as challenging for LLMs, especially for larger models such as GPT-4.
+We find that adaptation from English data is less effective, especially for Sundanese.
+However, GPT-4 Turbo can generate questions with some basic local knowledge in Indonesian and Sundanese when provided with human-created categories and concepts.
+  • We evaluate several LLMs on our dataset and find that they perform better on LLM-generated data than on human-generated data, indicating that the former is less challenging, especially for proprietary models like GPT-4.
+Nevertheless, many open-source LLMs still struggle to answer LLM-generated questions, highlighting significant room for improvement in these datasets.
 2 Related Work
 Commonsense Dataset
-Datasets and benchmarks for commonsense reasoning are crucial for developing models that understand and reason about real-world complexities.
+Datasets for commonsense reasoning are crucial for developing models that understand and reason about real-world complexities.
 Datasets like COPA Roemmele et al.
@@
 (2018), and Mcscript Ostermann et al.
-(2018) test commonsense reasoning via question-answering, but some questions require grade-school science knowledge.
+(2018) test commonsense reasoning via QA, but some questions require grade-school science knowledge.
 CommonsenseQA Talmor et al.
-(2019) presents the type of “purely” commonsense QA in the form of multiple-choice questions based on a concept knowledge from ConceptNet Speer et al.
+(2019) presents the type of “purely” commonsense QA in the form of multiple-choice questions based on knowledge from ConceptNet Speer et al.
 (2017), built through crowdsourcing.
-One question corresponds to one target concept, and other related concepts become distractors, acting as parts of the answer choices.
-However, since the dataset is crowdsourced, the crowdworkers bring their cultural background as part of their common knowledge, resulting in many questions that are “commonsense” only in the Western culture.
+However, the crowdworkers bring their cultural background as part of their common knowledge in building the data, resulting in many questions that are “commonsense” only in the Western culture.
 Cultural Evaluation Dataset
 Various datasets are used to assess language models’ ability to understand cultural nuances specific to a language.
-FORK Palta and Rudinger (2023) explores culinary cultural biases and assumptions.
-However, this dataset is relatively small, comprising only 184 questions and focusing solely on US, Indian, and Chinese customs.
-In the context of Indonesian, IndoMMLU Koto et al.
+FORK Palta and Rudinger (2023) explores culinary cultural biases and assumptions of US, Indian, and Chinese customs.
+IndoMMLU Koto et al.
 (2023) includes questions from Indonesian exams and covers regional cultural topics, such as the Minangkabau or Sundanese cultures.
 COPAL-ID Wibowo et al.
-(2023) is an Indonesian COPA-style dataset written by native speakers, thus incorporating more Indonesian cultural nuances compared to X-COPA.
+(2023) is a COPA-style dataset written by native speakers, thus incorporating more Indonesian cultural nuances compared to X-COPA.
 COPAL-ID is composed of approximately 500 questions, with a primary emphasis on the cultural aspects of the Jakarta region.
@@
 3.1 Commonsense QA: Definition and Scope
-The Cambridge Dictionary defines commonsense as "the basic level of practical knowledge and judgment that we all need to help us live in a reasonable and safe way." van Holthoon and Olson (1987) describe commonsense as "knowledge, judgment, and taste which is more or less universal and which is held more or less without reflection or argument." Both definitions highlight commonsense as the practical knowledge in daily life.
-Talmor et al.
-(2019) did not explicitly define the scope of commonsense in their QA dataset.
-Nevertheless, they outlined various skills required to answer the questions in their dataset correctly, such as spatial, cause & effect, social, activity, and many more.
-To build our dataset, we follow similar skills but focus on adding common or cultural knowledge specific to Indonesian and Sundanese contexts to reduce bias towards Western culture.
-Generally, one data point is a triple of concept, question, and options (with one correct answer).
-The data samples are shown in Figure 1.
+In our dataset, we focus on building questions that probe common or cultural knowledge in daily life within Indonesian and Sundanese contexts.
+Each data point is a triple of concept, question, and options, with one correct answer.
+To enhance the relevance of the dataset to local contexts and reduce Western cultural bias, we not only utilized existing question concepts from the English CommonsenseQA Talmor et al.
+(2019) but also manually created novel question concepts.
+We cover five categories: (1) culinary, (2) place, (3) culture, (4) history, and (5) activity, with 150 unique question concepts for each category, aiming for broad domain and knowledge coverage.
 3.2 Languages in Indonesia
-Indonesia is one of the most culturally and linguistically diverse countries in the world, with more than 700 languages spoken across the country Aji et al.
+Indonesia is one of the most culturally and linguistically diverse countries, with more than 700 languages spoken across the country Aji et al.
 (2022); Eberhard et al.
 (2021).
-Of the many languages in Indonesia, Indonesian is a unifying language used nationally.
-It utilizes the Latin script and was developed from literary "Classical Malay" of the Riau-Johor sultanate Sneddon (2003), with regional variations.
+Among the many languages, Indonesian is a unifying language used nationally.
+It utilizes the Latin script and was developed from literary "Classical Malay" Sneddon (2003), with regional variations.
 Over 80% of Standard Malay’s vocabulary is similar to Indonesian.
 Apart from Indonesian, regional languages like Sundanese are spoken by people of the same ethnicity.
-Sundanese is primarily spoken in West Java, as well as Banten, Jakarta, Central Java, and more.
-[^2] https://petabahasa.kemdikbud.go.id/infobahasa2.php?idb=56&idp=Jawa%20Barat Regional languages including Sundanese have influenced the formation and development of the Indonesian language.
-Both languages share similarities, such as their grammatical structure, but also differ significantly in aspects like the number of vowels and morphological features, including affixes.
+It is primarily spoken in West Java and is the second-largest regional language in Indonesia, with 34 million speakers Eberhard et al.
+(2021).
+Regional languages, including Sundanese, have influenced the formation and development of the Indonesian language.
+Both languages share similarities, such as their grammatical structure, but also differ significantly in aspects like the number of vowels (i.e., Sundanese has 2 additional vowels: é, eu) and morphological features, including affixes.
 We chose to study the national language and one regional language to illustrate the differences in the commonsense QA data generated via LLMs.
+It is also important to note that despite its large number of speakers, Sundanese has very limited data, and to date, there are no commonsense QA datasets available in Sundanese.
 4 Data Generation Methods
@@
 In addition, we employ humans to generate data for comparison.
-As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation, where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation, where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation, where we utilize LLMs to generate data based on the list of categories and concepts used in method (2).
+As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation (LLM_Adapt), where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation (Human_Gen), where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation (LLM_Gen), where we use LLMs to generate data based on the list of categories and concepts used in Human_Gen.
+As shown in Table 1, our final dataset contains 8,953 QA pairs in total (4,416 for Indonesian and 4,537 for Sundanese).
+Dataset Version Indonesian Sundanese
+Train / Valid / Test Total Train / Valid / Test Total
+LLM_Adapt 1,506 / 191 / 158 1,855 1,506 / 191 / 158 1,855
+Human_Gen 0 / 0 / 1,498 1,498 0 / 0 / 1,499 1,499
+LLM_Gen 0 / 0 / 1,063 1,063 0 / 0 / 1,183 1,183
+1,506 / 191 / 2,719 4,416 1,506 / 191 / 2,840 4,537
+Table 1: Statistics of our generated Indonesian and Sundanese CommonsenseQA dataset.
+While all prompts given to the LLM are in English, for the LLM_Adapt data, we specifically instruct the model to produce responses directly in Indonesian.
+Subsequently, we translate the resulting Indonesian data into Sundanese.
+In contrast, for the LLM_Gen data, we instruct the model to give the output directly in Indonesian and also Sundanese, to closely replicate the data generation process used for Human_Gen.
+The generation methods for each dataset are detailed below.
 4.1 Automatic Data Adaptation
@@
 (2019) to make it culturally relevant to Indonesian and Sundanese.
-We employ a multi-stage approach to adapt the data.
 First, we remove data that contains ambiguous options and offensive keywords.
@@
 We use GPT-3.5 Turbo results ensembled from five different prompts to determine the relevance of the data based on the three elements.
-Data that are considered irrelevant in at least one of these three elements are selected to be adapted.
+Data that are considered irrelevant in at least one of the three elements are selected to be adapted.
+The examples are shown in Table 11.
 The next step is to transform the selected irrelevant data.
@@
 The fully rephrased and adapted data are then translated from Indonesian to Sundanese using Google Translation API.
-Finally, we filter the data by examining the appearance of concepts in the questions and checking Sundanese-Indonesian back-translation with a similarity threshold of 90%.
+Finally, to ensure the dataset quality, we filter the data by examining the appearance of concepts in the questions and checking Sundanese-Indonesian back-translation with a similarity threshold of 90%.
 4.2 Manual Data Generation
-We employ annotators with different ethnicities from diverse regions covering Java and Bali island to ensure the variability of the knowledge in the human-generated dataset.
-The data collection process involves two main phases: (1) creating commonsense question-answer pairs and (2) answering commonsense questions.
-We perform quality control with additional human annotators to ensure the data quality.
+To build the human-generated dataset, we recruited 12 experienced annotators from diverse regions across Java and Bali, as detailed in Table 10.
+All are native speakers of Indonesian (and Sundanese), with a minimum of 15 years residing in their particular target language regions.
+Among them, 5 annotators have a degree in linguistics.
+These criteria ensure they possess both language proficiency and cultural familiarity.
+As for the data collection process, we have two main phases: (1) creating commonsense question-answer pairs and (2) answering commonsense questions.
+We also perform quality control to ensure the data quality.
+[^3] More details on annotators’ demographic information are presented in Appendix B.
+The complete annotation guideline is outlined in Appendix C.
 Creating Question-Answer Pairs
-We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept.
+We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept that we had newly created (§3.1).
 We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLM as much as possible.
-The categories and question concepts are manually curated to ensure wide domain coverage and to enrich the diversity of commonsense knowledge in the data.
-In total, we cover five categories: (1) culinary, (2) place, (3) culture, (4) history, and (5) activity, with 150 predefined question concepts for each category.
 Two annotators use the same 50 question concepts from each category to generate the questions, options, and answers.
@@
 After passing the quality assurance, the 1,500 QA pairs are then redistributed among all annotators for further review, where they are tasked with answering 1,250 commonsense questions each (excluding the set of data they have made in the first phase).
-This phase aims to evaluate the consistency and variance in the commonsense knowledge among annotators from various cultural backgrounds.
+Therefore, in this phase, each question is answered by five different annotators to capture the consistency and variance in the commonsense knowledge among annotators from various cultural backgrounds.
 Annotators are instructed not only to provide answers to the questions but also to comment on any ambiguities in the questions or options, or if they have any uncertainties when answering the questions.
 Quality Control
-We conduct Quality Control (QC) to ensure data correctness through manual human validation.
+We conduct Quality Control (QC) through manual validation to ensure data accuracy and maintain its “commonsense” nature.
 QC annotators manually review data for errors and provide feedback, which is then corrected by the annotator concerned.
 Once corrected, QC annotators re-check the data to determine whether it can be considered complete or still requires revision.
-An evaluation meeting with all annotators is also held to convey a more comprehensive evaluation regarding all types of errors and other findings, ensuring annotators have a better understanding and do not repeat similar errors.
+An evaluation meeting with all annotators is also held to convey a more comprehensive evaluation regarding all types of errors and other findings, ensuring annotators do not repeat similar errors and have a better understanding of the notion of commonsense.
 After the second phase (answering questions), we also excluded some questions if more than three annotators marked them as ambiguous.
@@
 To maintain dataset quality and uniqueness, we filter out duplicate entries and questions that do not explicitly contain the question concepts.
-Dataset Version Indonesian Sundanese
-Train Valid Test Train Valid Test
-LLM_Adapt 1,506 191 158 1,506 191 158
-Human_Gen - - 1,498 - - 1,499
-LLM_Gen - - 1,063 - - 1,183
-Total 1,506 191 2,719 1,506 191 2,840
-Table 1: Statistics of our generated Indonesian and Sundanese CommonsenseQA dataset.
-4.4 Resulting Data
-Our final CommonsenseQA dataset contains 4,416 QA pairs for Indonesian and 4,537 QA pairs for Sundanese.
-Table 1 provides detailed statistics for both languages.
-We also have three data variations: (1) LLM_Adapt: LLM-generated dataset constructed through data adaptation method (§4.1); (2) Human_Gen: Human-generated dataset, as explained in §4.2; (3) LLM_Gen: LLM-generated dataset given the same set of categories and concepts in Human_Gen4.3).
 5 Data Analysis
@@
 Concept Quality
-In the English (eng) to Indonesian (ind) concept adaptation, 95.57% are correctly adapted, including 16 out of 19 concepts needing major adaptation for cultural relevance, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall).
-Most adaptations are minor, primarily involving direct translations of eng to ind, as shown in Table 2.
-For Indonesian to Sundanese, the correct adaptation rate drops to 77.22%, reflecting weaker machine translation performance for Sundanese.
-The samples are shown in Table 3.
+In the English to Indonesian concept adaptation, 95.57% are correct, including 16 out of 19 concepts needing major adaptation for cultural relevance, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall), showing that the LLM has some basic knowledge to adapt English concepts, although most adaptations primarily involving exact translations without major alteration of the concept (Table 2).
+For Indonesian to Sundanese, the correct adaptation drops to 77.22%, reflecting weak machine translation performance in Sundanese (Table 8).
 Status Num (%) Concept Example
@@
 skate ice skating
-Table 2: Samples of English (eng) to Indonesian (ind) concept adaptation.
-Status Num (%) Concept Example
-Orig (ind) Modified (sun)
-Correct 122 (77.22%)
-hujan abu vulkanik
-(volcanic ashfall)
-hujan lebu vulkanik
-(volcanic ashfall)
-menara (tower) munara (tower)
-Wrong 36 (22.78%)
-cicak
-(house gecko)
-kadal imah
-(house lizard)
-klinik gigi
-(dental clinic)
-klinik dental
-(dental clinic)
-Table 3: Samples of Indonesian (ind) to Sundanese (sun) concept adaptation.
+Table 2: eng to ind concept adaptation result.
 Concept Variation
-Despite the high accuracy of concept adaptation, as shown in Figure 2, we observe that the adapted concepts are skewed towards ‘komodo’.
-This indicates that the model is biased toward a specific entity within a category, which could arise from its limited knowledge or the absence of a direct equivalent for the original English concept in Indonesian, leading it to default to one standard concept.
-This finding highlights the need to manually develop concepts across various categories to improve the dataset’s quality in terms of concept diversity and coverage.
+Despite the high accuracy of concept adaptation, as shown in Figure 2, the adapted concepts are skewed towards ‘komodo,’ indicating a bias towards a specific entity within a category.
+This could be due to the model’s limited knowledge or the absence of a direct equivalent for the original English concept in Indonesian, leading it to default to one standard concept.
+This finding highlights the limitations of adaptation and translation methods, and to improve dataset diversity and coverage, we need to manually develop concepts across various categories.
 Figure 2: Top-10 adapted question concepts taken from train, validation, and test set of LLM_Adapt data.
@@
 However, for Sundanese, the accuracy decreases significantly.
-In particular, the weak performance of the MT system in Sundanese is evident from its very low accuracy in LLM_Adapt.
-This suggests that synthetic data adaptation and generation is not useful for low-resource languages.
+In particular, the weak performance of the MT system in Sundanese is evident from its very low accuracy in LLM_Adapt, suggesting synthetic data adaptation/translation from English is not an optimal method for low-resource languages.
+In contrast, the LLM_Gen dataset, which involves generating synthetic data directly in the target language, shows a significantly higher number of correct questions, particularly for Sundanese.
+This indicates that direct generation in the target language is a more promising approach than adaptation or translation.
 Common Mistakes
-To further analyze the quality of the questions, we delve deeper into the mistakes that are commonly made, which are listed in Table 5.
+We manually reviewed the questions to identify common errors, detailed in Table 5.1.3.
 We observe that most of the errors in LLM_Adapt come from translation errors.
-For LLM_Gen, even though the accuracy of question generation is lower than LLM_Adapt, 54.26% of the errors in LLM_Gen are actually minor typos or mechanics errors.
-In the case of Sundanese, most errors in LLM_Gen are due to sentence fluency, indicating that the model is still struggling to generate fluent sentences in Sundanese.
+For Indonesian LLM_Gen, despite lower question generation accuracy, most of the errors are actually minor typos.
+In Sundanese LLM_Gen, however, the predominant issues relate to sentence fluency.
+This indicates that the model is still struggling to generate fluent sentences in Sundanese, where it occasionally uses phrases or words in Indonesian instead of Sundanese.
+While Indonesian and Sundanese share many linguistic features, they differ significantly in morphological features (§3.2).
+This characteristic, along with the smaller amount of Sundanese data used for training LLMs, results in the models being less exposed to Sundanese words, suffixes, or prefixes compared to Indonesian.
+This leads to the occasional appearance of Indonesian words in Sundanese data, even though the sentences may be grammatically correct.
 5.1.3 Options and Answer Analysis
-Options Quality
-We also evaluate the options in a similar manner to the evaluation of the question.
-From Table 6, we observe that LLM_Gen generates higher quality options compared to LLM_Adapt.
-However, there is still a significant gap in performance between Indonesian and Sundanese.
-This once again highlights the performance discrepancy between medium- and lower-resource languages.
-Common Mistakes
-We conduct a detailed analysis to identify the common errors found in each dataset, shown in Table 7.
-For LLM_Adapt, the most common issues are typos and mechanical errors, particularly with capitalization (e.g., "indonesia" instead of "Indonesia"), along with translation errors.
-In the Sundanese dataset, the errors are mainly due to major translation issues.
-As for LLM_Gen, although errors constitute less than 10% of the total evaluated samples, the primary issue is the presence of invalid options, particularly when no correct answers are among the options.
-For Sundanese, alongside translation errors, the model also produces some invalid options.
 Dataset % of correct questions
@@
 Table 3: Question generation accuracy of LLM-generated datasets. The accuracy reflects the percentage of a ‘clean’ question without any major or minor errors.
-Error Type % of error type
+Error Type % of questions
 LLM_Adapt LLM_Gen
 ind sun ind sun
-Translation 33.33% 48.51% 0.00% 13.61%
-Sent. structure 12.82% 11.94% 0.00% 0.00%
-Sent. fluency 28.21% 27.61% 36.17% 36.73%
-Sent. context 7.69% 2.24% 9.57% 16.33%
-Subjectivity 2.56% 0.75% 0.00% 0.00%
-Typo/mechanics 15.38% 8.96% 54.26% 33.33%
+No error 75.32% 15.19% 68.67% 51.00%
+[…]
+Sent. structure 3.16% 10.13% 0.00% 0.00%
+Sent. fluency 6.96% 23.42% 11.33% 18.00%
+Sent. context 1.90% 1.90% 3.00% 8.00%
+Subjectivity 0.63% 0.63% 0.00% 0.00%
+Typo/mechanics 3.80% 7.59% 17.00% 16.33%
 Table 4: Distribution of question generation error types of LLM-generated datasets.
-Dataset % of correct choices
+Dataset % of correct options
 ind sun
@@
 LLM_Gen 93.00% 58.67%
-Table 6: Choices generation accuracy of LLM-generated datasets. The accuracy reflects the percentage of ‘clean’ choices.
-Error Type % of error type
+Table 5: Options generation accuracy of LLM-generated datasets, measured by the % of clean/error-free options.
+Error Type % of options
 LLM_Adapt LLM_Gen
 ind sun ind sun
-Translation 10.17% 77.32% 0.00% 76.61%
-Sent. fluency 1.69% 1.03% 9.52% 3.23%
-Sent. context 1.69% 1.03% 0.00% 0.00%
-Invalid options 3.39% 2.06% 80.95% 20.16%
-Typo/mechanics 83.05% 18.56% 9.52% 0.00%
-Table 7: Distribution of choices generation error types of LLM-generated datasets.
-5.2 LLM vs. Human-Generated Data
-To compare the data generated by LLM with human-generated data, we analyze the proportion of shared unigram and bigram tokens between the LLM_Gen and Human_Gen datasets.
-The overlap percentage is calculated by dividing the number of shared tokens by the total number of tokens.
+No error 62.66% 38.61% 93.00% 58.67%
+[…]
+Sent. fluency 0.63% 0.63% 0.67% 1.33%
+Sent. context 0.63% 0.63% 0.00% 0.00%
+Invalid options 1.27% 1.27% 5.67% 8.33%
+Typo/mechanics 31.01% 11.39% 0.67% 0.00%
+Table 6: Distribution of options generation error types of LLM-generated datasets.
+Options Quality
+We also evaluate the options in a similar method to the evaluation of the question.
+From Table 5, we observe that LLM_Gen generates higher quality options compared to LLM_Adapt.
+However, there is still a significant gap in performance between Indonesian and Sundanese, once again highlighting the performance discrepancy between medium- and lower-resource languages.
+Common Mistakes
+As detailed in Table 5.1.3, for LLM_Adapt, the most common issues are minor typos and mechanical errors, particularly with capitalization (e.g., "indonesia" instead of "Indonesia").
+For Sundanese, the errors are mainly due to major translation errors.
+As for LLM_Gen, the most common errors involve the presence of invalid options, particularly when no correct answers are among the options.
+For Sundanese, alongside translation errors, the model also produces some invalid options.
+However, it is important to note that the proportion of invalid options is very low for both languages, indicating that the model is still capable of generating a large number of valid options.
+5.2 LLM vs. Human: Lexical Diversity
+To compare the lexical diversity between LLM and human-generated data, we analyze the proportion of shared tokens between the LLM_Gen and Human_Gen, calculated by dividing the number of unique shared tokens by the total number of unique tokens.
 We find that the unigram overlap percentage is 39.75% for Human_Gen and 65.48% for LLM_Gen.
@@
 This shows that many tokens present in Human_Gen also exist in LLM_Gen, but the reverse is not equally true.
-Given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol.
+We also find that Human_Gen dataset has 8,596 unique tokens, higher than LLM_Gen with 6,677 unique tokens.
+Upon some sample-level analysis, we find that given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol.
 [^4] In the context of culinary, ‘kalis’ means a state of dough that is well-kneaded and ready to be processed further.
 ‘Cimol’ is an Indonesian street food made from tapioca flour.
-Upon sample-level analysis, we also find token variance disparities, particularly related to cultural context.
+Although LLM can still generate some valid concepts, it tends to ask more general questions or popular concepts.
 For instance, given ‘kerupuk’ (crackers) concept, human annotators can formulate questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers).
-In contrast, LLM tends to ask more general questions, like "What is the common primary ingredient of crackers?" Also, although LLM can generate some culturally related entities such as ‘tari kecak’, a traditional dance from Bali, it is more popular compared to ‘tari pendet’, which is generated by the human annotator.
-More examples are outlined in Table 11.
+In contrast, LLM tends to ask more general (but still relevant) questions, like "What is the common primary ingredient of crackers?" More examples are shown in Table 12.
 6 Benchmark Result
@@
 We include PolyLM-13B Wei et al.
-(2023), an LLM trained on a multilingual dataset, predominantly in English and Chinese; BLOOMZ-7B Muennighoff et al.
-(2022), an LLM fine-tuned on the xP3 dataset; and SeaLLM-7B Nguyen et al.
-(2023), an LLM with a focus on Southeast Asian languages.
+(2023), which trained on a multilingual dataset (mostly English and Chinese); BLOOMZ-7B Muennighoff et al.
+(2022), which fine-tuned on xP3; SeaLLM-7B Nguyen et al.
+(2023), which covers Southeast Asian languages; and Aya Üstün et al.
+(2024), recent open LLM trained on 101 languages, including Indonesian and Sundanese.
 Monolingual LLMs
-We include Merak-v4 Ichsan (2023), an Indonesian LLM fine-tuned based on MistralOrca-7B. We also include MalaysianMistral-7B Zolkepli et al. (2024), a Malaysian LLM extended from Mistral-7B.
+We include Merak-v4 Ichsan (2023), an Indonesian LLM based on MistralOrca and MalayMistral-7B Zolkepli et al. (2024), a Malaysian LLM extended from Mistral.
 Proprietary LLMs
-We include GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo, OpenAI (2023), which are accessed via OpenAI API.
+We include GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo OpenAI (2023).
 We use three prompt variations for all models (see Appendix E).
@@
 We use accuracy as the evaluation metric.
+Figure 3: Overall performance of LLMs on our combined test set.
+Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen dataset in Indonesian and Sundanese.
 6.2 Overall Performance
-Figure 3: Overall performance of LLMs on our combined test set. We run three different prompts for each question in a zero-shot setting.
 […]
+We also observe a substantial gap between Indonesian and Sundanese (10–20% accuracy drop), suggesting that current LLMs struggle to answer our Sundanese commonsense questions, even in multiple-choice settings.
+The huge gap, particularly in Merak-v4, highlights the limitations of training LLMs solely on Indonesian texts, as it does not ensure transferable performance across other local languages, especially due to the morphological differences between the languages.
+Including Sundanese texts in the training data, as demonstrated by improved performance in the Aya model, has proven effective in narrowing the performance gap between Indonesian and Sundanese.
 6.3 LLM vs. Human-Generated Data
-Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen.
-To assess LLMs’ ability to answer both LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets.
-Our result, shown in Figure 4, reveals that LLMs perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data.
-This indicates that the model may possess some knowledge to answer LLM-generated questions, even though it was not specifically trained on Indonesian or Sundanese data.
-However, the model still struggles to answer new knowledge present in Human_Gen data.
-The performance gap is also observed in Indonesian LLM, Merak-v4.
-The accuracy of this model in Sundanese is also low, indicating that training LLM solely on Indonesian texts is not enough to make it perform well in local languages such as Sundanese.
+To assess LLMs’ ability to answer LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets.
+Figure 4 shows that LLMs generally perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data.
+This indicates that the model may possess some knowledge to answer LLM_Gen questions, even though it was not specifically trained on Indonesian or Sundanese data.
+Still, it is important to note that even in LLM_Gen, MistalOrca’s average score is still lower compared to LLMs that include Indonesian texts in their training.
+Interestingly, as the model gets stronger, we observe that the gap tends to narrow, especially in GPT models.
+For Aya, despite its relatively stronger performance compared to other open-source LLMs, its performance is significantly dropped in Human_Gen data.
+This decline is likely due to Aya being a ‘translation-heavy’ model, with a large proportion of its training data derived from translations rather than human annotations.
 6.4 Performance by Question Category
-Figure 5: LLMs performance by question category in LLM_Gen and Human_Gen for Indonesian and Sundanese.
 Figure 8 shows the LLMs’ performance across different question categories.
-It reveals that LLMs perform better in the activity and place questions, but struggle in culinary.
-A notable performance gap exists between LLM_Gen and Human_Gen dataset, with weaker results on the latter.
-Moreover, open LLMs exhibit a significant performance gap compared to proprietary models, especially in Human_Gen, indicating their shortcomings in capturing commonsense knowledge.
-Interestingly, a significant improvement is seen from GPT-4 (gpt-4-0613) to GPT-4 Turbo (gpt-4-1106), especially in Sundanese Human_Gen.
-This suggests that the model is acquiring more knowledge, one reason for which possibly comes from the interactions on the ChatGPT web.
-Despite these advancements, some categories still score below 80%, showing there is still room for improvement.
+It shows that LLMs perform better in the activity and place questions, which benefit from a lot of information readily available on the internet.
+However, they struggle with culinary questions, where specialized or cultural knowledge is often required.
+Interestingly, a significant improvement is seen from GPT-4 (0613 ver.) to GPT-4 Turbo (1106 ver.), particularly in Sundanese Human_Gen.
+This suggests that the model is acquiring more knowledge, which possibly comes from the interactions on the ChatGPT web.
+However, some categories still score below 80%, showing there is still room for improvement.
 7 Discussion
 7.1 Multiple-Choice vs. ‘Free’ Generation
-Since our dataset is in a multiple-choice format, LLMs might look better than they actually are because they can just pick one answer from the given options.
-To see LLMs’ genuine capability, we test GPT-4 Turbo on 100 randomly sampled questions from Indonesian Human_Gen in open-ended settings, i.e., asking the question directly without showing answer options or providing any extra instructions.
-Our manual evaluation shows a 77% accuracy rate, highlighting the model’s struggle with open-ended questions and confirming our dataset’s importance.
-Among the wrong answers, 18 cases stemmed from overly general answers not specific to Indonesia.
-For example, when asked "Where can we find many villas?" the model answers "Bali" but also listed other cities in mostly Western countries.
-In contrast, the annotator writes "Bandung" as the correct answer.
-As for the other wrong cases, the model struggles the most in answering questions in the culinary category.
+[…]
+Among the wrong answers, most cases come from overly general answers not specific to Indonesia, especially for questions in culinary, culture, and activity categories.
+For instance, given unique activity questions such as: "What is the mandatory song to be sung during the ‘mengheningkan cipta’ (moment of silence) in the flag ceremony?" the model gives a wrong answer in open-ended generation settings even if it could identify the correct answer in the multiple-choice setup.
 More failure examples can be seen in Table 13.
-7.2 Extension to Other Local Languages
+7.2 LLM-generated Data Quality
+Due to the high performance of LLMs in answering LLM-generated data, one might argue that the data is less valuable.
+Ideally, creating a larger human-generated dataset is the most straightforward method to create a high-quality dataset.
+However, this approach may not always be feasible in the context of underrepresented languages, particularly given the limited resources.
+Building a dataset manually can be costly and challenging (we spent more than \$3,000 in total to build Human_Gen), and leveraging LLMs could be a practical and cheaper alternative, especially if we want to scale up the data.
+To address the limitation of LLM-generated data, we can apply additional processes to reduce the noise.
+For instance, a collaboration between humans and models can be conducted to fix and revise the synthetic questions Liu et al.
+(2022); Putri and Oh (2022).
+This noise can also be reduced automatically, especially in Indonesian LLM_Gen questions, where errors typically involve minor typos.
+It is also important to note that not all LLM-generated data have low quality.
+Our experiments demonstrate that LLM_Gen has significantly fewer errors compared to LLM_Adapt, particularly in Sundanese.
+We also compared the performance between the cleaned vs.
+raw version of this data (Appendix F) and found that the performance variance between the original and cleaned data of LLM_Gen is relatively minor.
+8 Conclusion
+In this study, we introduced new Indonesian and Sundanese CommonsenseQA datasets built using various methods, including automatic dataset adaptation and generation with LLMs, as well as manual dataset creation by human annotators from diverse regions.
+Our thorough analysis reveals that adaptation from the existing English data is less effective for Sundanese.
+In contrast, when directly generated in the target languages, GPT-4 Turbo can produce relevant knowledge in both languages, although the cultural depth has not yet matched the human-generated ones.
+Nevertheless, given the limited resources, combining LLMs and humans for dataset creation can be a practical solution, allowing for a more efficient dataset creation by reducing the over-reliance on a single data source.
+Our goal is for LLMs to serve as beneficial tools for diverse communities, not solely those from higher-resource languages, cultures, or regions.
+This is especially important since LLMs are continuously being used to generate data.
+We hope our work represents a crucial step toward achieving this broader objective.
+Limitations
+Language and Region Coverage
+In terms of language coverage, we were only able to cover Indonesian and Sundanese due to the available resources and the authors’ familiarity with these languages.
+Additionally, the annotators we recruited were mostly from Java island, with one annotator from Bali island.
+Despite our effort to include a range of question concepts from different regions, including those beyond Java and Bali islands, it is possible that some bias may exist, especially in the Indonesian dataset.
+This is because the questions were generated primarily by annotators from Java and Bali, and their perspectives and cultural backgrounds may have influenced the content.
+Nonetheless, we have taken measures to eliminate potentially harmful or stereotypical questions.
+Extension to Other Local Languages
 Besides Indonesian, our study focuses on one Indonesian local language, Sundanese.
@@
 (2023).
-Generating good-quality synthetic data could be particularly beneficial for extremely low-resource languages, such as Buginese or Toba Batak, given the difficulty of finding native speakers.
-Yet, our case study in Sundanese indicates that LLMs might struggle even more with understanding cultural nuances in such languages.
-8 Conclusion
-In this study, we created Indonesian and Sundanese CommonsenseQA datasets using various dataset collection methods, including automatic dataset generation utilizing LLMs and manual dataset generation employing human annotators from diverse regions.
-Our thorough analysis reveals a substantial gap in the performance of LLMs between Indonesian and Sundanese, both in the dataset generation and in answering the questions in the dataset itself.
-Thus, one must be cautious when using existing LLMs in generating datasets, particularly for lower-resource languages, since their culturally relevant commonsense knowledge may be limited.
-Our findings highlight the urgent need to create high-quality resources for these underrepresented languages to prevent them from being left behind.
-Limitations
-Our work has some limitations that should be acknowledged.
-Firstly, in terms of language coverage, we were only able to cover Indonesian and Sundanese due to the available resources and the authors’ familiarity with these languages.
-Additionally, the annotators we recruited were mostly from Java island, with one annotator from Bali island.
-Despite our effort to include a range of question concepts from different regions, including those beyond Java and Bali islands, it is possible that some bias may exist, especially in the Indonesian dataset.
-This is because the questions were generated primarily by annotators from Java and Bali, and their perspectives and cultural backgrounds may have influenced the content.
-Nonetheless, we have taken measures to eliminate potentially harmful questions.
-Additionally, we use a multiple choice question format, following English CommonsenseQA data format Talmor et al.
+Generating synthetic data could be particularly beneficial for extremely low-resource languages, given the difficulty of finding native speakers.
+Yet, our case study in Sundanese indicates that LLMs might struggle even more with generating fluent sentences and/or culturally unique concepts in such languages.
+Still, we believe that the combination or even collaboration between LLM and humans may serve as a starting point.
+Multiple-Choice Format
+We use a multiple choice question format, following English CommonsenseQA data format Talmor et al.
 (2019) to facilitate a more straightforward and robust evaluation process.
-Although open-ended generation may offer a more challenging benchmark for LLMs, as discussed in Section 7.1, evaluating LLMs in such settings poses its own set of challenges, especially in low-resource languages where ‘LLM-as-a-judge’ approach may not be as effective as in English.
+Although open-ended generation may offer a more challenging benchmark for LLMs, as also discussed in Section 7.1, evaluating LLMs in such settings poses its own set of challenges, especially in low-resource languages where ‘LLM-as-a-judge’ approach may not be as effective as in English.
 Nevertheless, our dataset can serve as a starting point for this line of research direction.
@@
 Table 7: Win rate comparison of Merak-v4 (open Indonesian LLM) and GPT-4 Turbo (best-performing proprietary LLM).
-Table 8 shows the win rate accuracy of concept, question, and choices generation of Indonesian LLM, Merak-v4, and best-performing proprietary LLM, GPT-4 Turbo.
+Table 7 shows the win rate accuracy of concept, question, and options generation of Indonesian LLM, Merak-v4, and best-performing proprietary LLM, GPT-4 Turbo.
 The results demonstrate that GPT-4 Turbo significantly outperforms Merak-v4, with win rates ranging from 72% to 96%.
@@
 For instance, it generates "Siapa yang memimpin kampanye pemilihan presiden pertama di Amerika Serikat?" (eng: Who led the first presidential election campaign in the United States?).
-A.3 Additional Analysis of the Generated Data: Common Mistakes in Concept Adaptation
+A.3 Additional Analysis of the LLM-generated Data: Common Mistakes in Concept Adaptation
+Status Num (%) Concept Example
+Orig (ind) Modified (sun)
+Correct 122 (77.22%)
+hujan abu vulkanik
+(volcanic ashfall)
+hujan lebu vulkanik
+(volcanic ashfall)
+menara (tower) munara (tower)
+Wrong 36 (22.78%)
+cicak
+(house gecko)
+kadal imah
+(house lizard)
+klinik gigi
+(dental clinic)
+klinik dental
+(dental clinic)
+Table 8: Samples of Indonesian (ind) to Sundanese (sun) concept adaptation.
 Error Type Num (%) of errors
-eng normal-$\rightarrow$ ind ind normal-$\rightarrow$ sun
+eng $\rightarrow$ ind ind $\rightarrow$ sun
 Translation 4 (57.14%) 31 (86.11%)
@@
 In the case of adapting concepts from ind to sun, similar to eng to ind, the majority of errors also arise from translation errors, with a larger number of errors.
-Appendix B Human Annotation Guideline
+Appendix B Annotators Demographics
+Data Team Ethnicity Domicile Num
+Indonesian Sundanese Sukabumi, West Java 1
+Depok, West Java 1
+Javanese Magelang, Central Java 1
+Bojonegoro, East Java 1
+Betawi Tangerang, Banten 1
+Balinese Denpasar, Bali 1
+Sundanese Sundanese Bandung, West Java 2
+Bogor, West Java 1
+Majalengka, West Java 1
+Sukabumi, West Java 2
+Total 12
+Table 10: Demographic information of the annotators from each dataset team.
+Note that even though Depok is included in West Java Province and Tangerang is included in Banten Province, both are geographically closer to Jakarta and considered as part of the Greater Jakarta area (Jabodetabek).
+In accordance with the aim of constructing the dataset, we involved 6 Indonesian and 6 Sundanese native-speaker annotators.
+We chose annotators who have prior experience in building NLP datasets in Indonesian and/or Sundanese through close recruitment.
+All the annotators who worked on the Indonesian dataset were people from several regions on the islands of Java and Bali.
+Meanwhile, the annotators involved in building the Sundanese dataset were Sundanese people who come from several different regions in West Java.
+The detailed annotators’ demographics are shown in Table 10.
+Appendix C Human Annotation Guideline
 To ensure a high-quality and standardized format for question-answers annotation, we provide a specific guideline during the annotation process.
@@
 The total number of concepts that will be accepted is 250 concepts, equivalent to the workload for each person, where the expectation is to create one question per concept.
-Appendix C Annotators Demographics
-Data Team Ethnicity Domicile Num
-Indonesian Sundanese Sukabumi, West Java 1
-Depok, West Java 1
-Javanese Magelang, Central Java 1
-Bojonegoro, East Java 1
-Betawi Tangerang, Banten 1
-Balinese Denpasar, Bali 1
-Sundanese Sundanese Bandung, West Java 2
-Bogor, West Java 1
-Majalengka, West Java 1
-Sukabumi, West Java 2
-Total 12
-Table 10: Demographic information of the annotators from each dataset team.
-Note that even though Depok is included in West Java Province and Tangerang is included in Banten Province, both are geographically closer to Jakarta and considered as part of the Greater Jakarta area (Jabodetabek).
-In accordance with the aim of constructing the dataset, we involved all Indonesian native-speaker annotators in the manual data generation process.
-All the annotators who worked on the Indonesian dataset were people from several regions on the islands of Java and Bali.
-Meanwhile, the annotators involved in building the Sundanese dataset were Sundanese people who come from several different regions in West Java.
-Each dataset development team consists of 6 annotators.
-The detailed annotators’ demographics are shown in Table 10.
 Appendix D Manual Data Generation Findings
 During the manual data generation process, we discovered several interesting findings from the data created by the annotators.
+Note that the common errors within these findings occur before the quality control meeting (§4.2).
+In our final dataset, we have ensured that such problems have been fixed.
 Questions
 We found quite a lot of interesting findings in the process of creating commonsense questions.
-We found that there were questions created using very general contexts, while the limitations of this data generation were the Indonesian and Sundanese contexts (specifically for the Sundanese dataset).
-For instance, "Apa jenis restoran yang paling terkenal di seluruh dunia?" (eng: Which type of restaurant is most famous worldwide?) In that question, the annotator developed the concept question “restoran” (eng: restaurant) into a question with a very general context: “di seluruh dunia" (eng: all over the world).
+We found that there were questions created using very general contexts, while we required the data generation to be scoped within the Indonesian and Sundanese contexts (specifically for the Sundanese dataset).
+For instance, an annotator produced "Apa jenis restoran yang paling terkenal di seluruh dunia?" (eng: Which type of restaurant is most famous worldwide?) In that question, the annotator developed the concept question “restoran” (eng: restaurant) into a question with a very general context: “di seluruh dunia" (eng: all over the world).
 We also found subjective questions, so the answers that emerged also had subjective value (not commonsense).
@@
 This seems to be based on confusion from annotators regarding the boundaries of commonsense and logic/causation.
-Choices
-The findings related to choices that are most often encountered are making choices that have the same value as each other so that it is difficult to determine the answer and distractor.
+Options
+The findings related to options that are most often encountered are making choices that have the same value as each other so that it is difficult to determine the answer and distractor.
 For instance, the question "Kue apa yang biasanya disajikan pada momen lebaran?" (eng: What cakes are usually served during Eid?).
@@
 Answers and distractors are another category that also has many findings in the manual data generation process.
-This issue is not much different from the findings in Choices: answers and distractors are equivalent.
+This issue is not much different from the findings in Options: answers and distractors are equivalent.
 Apart from that, in this category, it was also found that personal experience was used in determining answers so that more common answers were determined as distractors.
@@
 We examine this by calculating the number of answer conflicts that arise during the "answering question" phase of our data generation pipeline (§4.2).
-The results from Indonesian and Sundanese annotators are detailed in Figure 6 and 7, respectively.
+The results from Indonesian and Sundanese annotators are detailed in Figure 5 and 6, respectively.
 Our analysis revealed that, out of the Indonesian annotators, the one from Bali has the highest number of conflicting answers.
@@
 However, caution is still necessary as errors can still occur.
+Classification Original Data (eng) LLM Adapted Data (ind)
+concept=irrelevant
+name=irrelevant
+option=irrelevant
+Concept: koala
+Question: James looked for koalas, but
+misunderstood and went to the wrong environment
+where is it impossible to find koalas?
+Options:
+A. jungle
+B. great outdoors
+C. siberia
+D. queensland
+E. wilderness
+Concept: orangutan
+Question: Budi mencari orangutan, tetapi
+salah paham dan pergi ke lingkungan yang salah.
+Di mana mustahil untuk menemukan orangutan?
+Options:
+A. hutan hujan tropis
+B. alam bebas
+C. gurun sahara
+D. kalimantan
+E. belantara
+concept=relevant
+name=irrelevant
+option=irrelevant
+Concept: hotel room
+Question: John rented a hotel room while on a
+business trip. It was a company sponsored event.
+Where might it have been?
+Options:
+A. convention center
+B. business interviews
+C. michigan
+D. resort
+E. city
+Concept: kamar hotel
+Question: Andi menyewa kamar hotel saat dalam
+perjalanan bisnis. Ini adalah acara yang disponsori
+perusahaannya. Di mana acara tersebut mungkin
+diadakan?
+Options:
+A. pusat konvensi
+B. wawancara bisnis
+C. jakarta
+D. resor
+E. kota
+concept=irrelevant
+name=relevant
+option=irrelevant
+Concept: blue ridge mountains
+Question: Which state is the blue ridge mountains
+roanoke city in?
+Options:
+A. appalachians
+B. north carolina
+C. virginia
+D. pennsylvania
+E. montana
+Concept: Gunung Bromo
+Question: Di provinsi mana Gunung Bromo terletak?
+Options:
+A. jawa barat
+B. jawa tengah
+C. jawa timur
+D. yogyakarta
+E. bali
+concept=relevant
+name=relevant
+option=relevant
+Concept: teacher
+Question: What does a kindergarten teacher do
+before nap time?
+Options:
+A. lower expectations
+B. encourage
+C. fear
+D. time test
+E. tell story
+N/A
+Table 11: Examples of the concept, name, and option relevancy classification result. The English data instance that is classified as irrelevant in any of these three elements will be adapted.
 Category Concept Data Creator Question Choices Answer
@@
 We provide some GPT-4 Turbo generation samples when given the question in ‘free’ or open-ended generation settings in Table 13.
+Figure 8: LLMs performance by question category in LLM_Gen and Human_Gen for Indonesian and Sundanese.

### v3

@@
 It is not yet clear whether current LLMs can create QA data that adequately includes the cultural nuances specific to certain languages.
-Another common method for constructing non-English datasets is by utilizing a machine translation model.
+Another common method for constructing non-English datasets is by using machine translation.
 Although more scalable, this method cannot be straightforwardly applied due to the contextual irrelevancy of the data, primarily due to geographical differences (which could also influence cultural differences).
@@
 Each data point is a triple of concept, question, and options, with one correct answer.
-To enhance the relevance of the dataset to local contexts and reduce Western cultural bias, we not only utilized existing question concepts from the English CommonsenseQA Talmor et al.
-(2019) but also manually created novel question concepts.
-We cover five categories: (1) culinary, (2) place, (3) culture, (4) history, and (5) activity, with 150 unique question concepts for each category, aiming for broad domain and knowledge coverage.
+To enhance the relevance of the dataset to local contexts and reduce Western cultural bias, we incorporate both existing question concepts from the English CommonsenseQA Talmor et al.
+(2019) and novel question concepts that we manually created.
+Unlike English CommonsenseQA, our dataset includes question category metadata, covering five categories: culinary, place, culture, history, and activity.
+Each category has 150 unique question concepts, ensuring broad domain and knowledge coverage.
+[^2] More details on how we curate our categories and concepts can be seen in Appendix B.
+Another key difference is that English CommonsenseQA concepts are sourced from a knowledge base, and they divide the concept sets into questions and options.
+In contrast, our dataset provides only a question concept and its category as input.
+The options, including the correct answer, are entirely generated by the data creator.
 3.2 Languages in Indonesia
@@
 4 Data Generation Methods
-To investigate whether LLMs can generate culturally relevant commonsense QA data in Indonesian and Sundanese, we construct a dataset using various methods with LLMs as data generators.
-In addition, we employ humans to generate data for comparison.
+To investigate whether LLMs can generate culturally relevant commonsense QA data in Indonesian and Sundanese, we build a dataset using various methods with LLMs as data generators.
+We also employ humans to generate data for comparison.
 As illustrated in Figure 1, we apply three dataset generation methods: (1) Automatic Data Adaptation (LLM_Adapt), where we leverage LLMs to automatically adapt English CommonsenseQA data to our target languages; (2) Manual Data Generation (Human_Gen), where we ask native-speaker human annotators to manually construct the dataset; and (3) Automatic Data Generation (LLM_Gen), where we use LLMs to generate data based on the list of categories and concepts used in Human_Gen.
-As shown in Table 1, our final dataset contains 8,953 QA pairs in total (4,416 for Indonesian and 4,537 for Sundanese).
+Table 1 shows the statistics of our final dataset, containing a total of 8,953 QA pairs (4,416 Indonesian and 4,537 Sundanese).
 Dataset Version Indonesian Sundanese
@@
 1,506 / 191 / 2,719 4,416 1,506 / 191 / 2,840 4,537
-Table 1: Statistics of our generated Indonesian and Sundanese CommonsenseQA dataset.
+Table 1: Statistics of our generated Indonesian and Sundanese CommonsenseQA dataset. We retained the original English CommonsenseQA splits in LLM_Adapt to avoid data contamination.
 While all prompts given to the LLM are in English, for the LLM_Adapt data, we specifically instruct the model to produce responses directly in Indonesian.
@@
 4.1 Automatic Data Adaptation
-We build the first LLM-generated data by adapting the CommonsenseQA dataset Talmor et al.
-(2019) to make it culturally relevant to Indonesian and Sundanese.
-First, we remove data that contains ambiguous options and offensive keywords.
-Next step involves assessing three elements: concept, name, and location, in order to sample the data that will be adapted.
-We take the question concepts from the existing CommonsenseQA data.
-As for the name and location extraction, we use Stanford’s Stanza NLP library Qi et al.
-(2020) and ConceptNet Speer et al.
-(2017) API.
-We use GPT-3.5 Turbo results ensembled from five different prompts to determine the relevance of the data based on the three elements.
+We build the first LLM-generated data by adapting English CommonsenseQA dataset Talmor et al. (2019) to make it culturally relevant to Indonesian and Sundanese.
+Data Selection
+To select the data to be adapted, we first remove data containing offensive keywords and those with duplicate or similar options in different tenses.
+Unlike English, Indonesian and Sundanese do not change form to indicate tenses (e.g., ‘passed’ and ‘passing’ both translate to ‘lulus’).
+Then, we sample the data by assessing three elements: concept, name, and location.
 Data that are considered irrelevant in at least one of the three elements are selected to be adapted.
-The examples are shown in Table 11.
+We take question concepts from the existing CommonsenseQA data and use Stanza Qi et al.
+(2020) and ConceptNet API [^3] https://github.com/commonsense/conceptnet5/wiki/API for name and location extraction.
+[…]
+QA Pairs Adaptation
 The next step is to transform the selected irrelevant data.
@@
 to rephrase the sampled questions and options to align with Indonesian cultures.
-Subsequently, for data flagged with Western-centric names, a second rephrasing is conducted using the GPT-3.5 Turbo to replace all person names.
-The fully rephrased and adapted data are then translated from Indonesian to Sundanese using Google Translation API.
-Finally, to ensure the dataset quality, we filter the data by examining the appearance of concepts in the questions and checking Sundanese-Indonesian back-translation with a similarity threshold of 90%.
+Subsequently, for data flagged with Western-centric names, we use GPT-3.5 Turbo to replace all person names.
+The fully adapted data are then translated from Indonesian to Sundanese using Google Translation API.
+We choose machine translation due to the unreliability of direct adaptation, as the concepts are sourced from English data.
+Our preliminary run on eng to sun adaptation shows that 90% of the samples contained errors, with half due to hallucinations.
+This occurred less frequently in eng to ind adaptation.
+For example, for the ‘bald eagle’ concept, GPT-4 Turbo generated ‘elang Jawa’ (Javan hawk-eagle) in Indonesian, but it generated the non-existent concept ‘garuda puspa’ (literally translates to eagle flower) in Sundanese.
+Although not ideal, translation from the adapted Indonesian QA pairs generally provides better data quality.
+Data Filtering
+Finally, we eliminate low-quality data by removing instances where the concepts do not appear in the questions.
+We also utilize a back-translation method to filter out poor translations of Indonesian (ind) to Sundanese (sun) data.
+We discard QA pairs if the similarity between the original Indonesian (ind) and the back-translated version (ind’) is less than 0.9.
+We measure similarity using an embedding-based metric with multilingual MiniLM Reimers and Gurevych (2019) and LaBSE Feng et al.
+(2022) models, and keep the data if either similarity score is above 0.9.
 4.2 Manual Data Generation
-To build the human-generated dataset, we recruited 12 experienced annotators from diverse regions across Java and Bali, as detailed in Table 10.
+To build the human-generated dataset, we recruited 12 experienced annotators from diverse regions across Java and Bali.
 All are native speakers of Indonesian (and Sundanese), with a minimum of 15 years residing in their particular target language regions.
@@
 We also perform quality control to ensure the data quality.
-[^3] More details on annotators’ demographic information are presented in Appendix B.
-The complete annotation guideline is outlined in Appendix C.
-Creating Question-Answer Pairs
+[^6] More details on annotators’ demographic information and annotation guidelines are presented in Appendix C and D.
+QA Pairs Creation
 We first instruct the annotators to create commonsense question-answer pairs based on the given category and question concept that we had newly created (§3.1).
-We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLM as much as possible.
-Two annotators use the same 50 question concepts from each category to generate the questions, options, and answers.
-This process produces 1,500 QA pairs for each language in total.
+We also ask them to rely on their existing knowledge when making questions and avoid using internet search or LLMs as much as possible.
+Each concept is annotated by 2 annotators, with each annotator covering all 5 categories to ensure consistency.
+Since we have 150 unique concepts per category and 6 annotators per language, each annotator is assigned to create QA pairs for 50 concepts per category.
+This results in 300 QA pairs per category (150 concepts $\times$× 2 annotators) and a total of 1,500 QA pairs per language (300 QA pairs $\times$× 5 categories).
 Answering Question
@@
 We adopt a batching approach to streamline the data generation process, providing the model with a maximum of 5 distinct question concepts from the same categories in one API call.
+We chose 5 concepts because this number is optimal; larger batches would exceed the API’s maximum length, while smaller batches would result in longer processing times.
 To maintain dataset quality and uniqueness, we filter out duplicate entries and questions that do not explicitly contain the question concepts.
@@
 5.1 LLM-Generated Data
-To evaluate the quality of our LLM-generated data, we manually reviewed all 158 samples from the LLM_Adapt test set.
+To evaluate the quality of our LLM-generated data, we manually reviewed all 158 samples from the LLM_Adapt test set (see Table 1).
 Based on the evaluation, we then calculate the accuracy of the generated concepts, questions, and options.
-Furthermore, we evaluate 300 randomly selected samples from the LLM_Gen using the same procedure.
-5.1.1 Concept Analysis
-Concept Quality
-In the English to Indonesian concept adaptation, 95.57% are correct, including 16 out of 19 concepts needing major adaptation for cultural relevance, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall), showing that the LLM has some basic knowledge to adapt English concepts, although most adaptations primarily involving exact translations without major alteration of the concept (Table 2).
-For Indonesian to Sundanese, the correct adaptation drops to 77.22%, reflecting weak machine translation performance in Sundanese (Table 8).
+We also evaluate 300 randomly selected samples from the LLM_Gen using the same procedure.
 Status Num (%) Concept Example
@@
 Table 2: eng to ind concept adaptation result.
+Status Num (%) Concept Example
+Orig (ind) Modified (sun)
+Correct 122 (77.22%)
+hujan abu vulkanik
+(volcanic ashfall)
+hujan lebu vulkanik
+(volcanic ashfall)
+menara (tower) munara (tower)
+Wrong 36 (22.78%)
+cicak
+(house gecko)
+kadal imah
+(house lizard)
+klinik gigi
+(dental clinic)
+klinik dental
+(dental clinic)
+Table 3: ind to sun concept adaptation result.
+5.1.1 Concept Analysis
+Concept Quality
+To evaluate the quality of the adapted concepts, we regard the concept as correct if it is a real, existing concept (not a hallucination) and relevant to the Indonesian or Sundanese context.
+As shown in Table 2, 95.57% of English to Indonesian adaptations are correct, including 16 out of 19 concepts needing major adaptation, such as ‘snow’ to ‘hujan abu vulkanik’ (volcanic ashfall).
+This result indicates that the LLM can adequately adapt some English concepts to Indonesian, although most involve direct translations without major alterations.
+However, for Indonesian to Sundanese (Table 3), the correct adaptation drops to 77.22%, reflecting weak machine translation (MT) performance in Sundanese.
 Concept Variation
 Despite the high accuracy of concept adaptation, as shown in Figure 2, the adapted concepts are skewed towards ‘komodo,’ indicating a bias towards a specific entity within a category.
-This could be due to the model’s limited knowledge or the absence of a direct equivalent for the original English concept in Indonesian, leading it to default to one standard concept.
-This finding highlights the limitations of adaptation and translation methods, and to improve dataset diversity and coverage, we need to manually develop concepts across various categories.
+This could be due to the model being trained on data with insufficient knowledge or the absence of a direct equivalent of the English concept in Indonesian, leading it to default to one standard concept.
+This finding highlights the limitations of concept adaptation from existing English data.
+To improve dataset diversity and coverage, manual concept creation in the target language is needed.
 Figure 2: Top-10 adapted question concepts taken from train, validation, and test set of LLM_Adapt data.
@@
 To evaluate the quality of the generated questions, we apply a strict criterion: any errors, even minor, are marked as incorrect.
-As shown in Table 3, the Indonesian datasets show a high percentage of error-free questions, ranging between 68–75%.
+As shown in Table 4, the Indonesian datasets have a high percentage of error-free questions, ranging between 68–75%.
 However, for Sundanese, the accuracy decreases significantly.
@@
 For Indonesian LLM_Gen, despite lower question generation accuracy, most of the errors are actually minor typos.
-In Sundanese LLM_Gen, however, the predominant issues relate to sentence fluency.
-This indicates that the model is still struggling to generate fluent sentences in Sundanese, where it occasionally uses phrases or words in Indonesian instead of Sundanese.
+In Sundanese LLM_Gen, however, the predominant issues relate to sentence fluency, indicating ongoing challenges in automatically generating smooth and fluent sentences in Sundanese.
+We also found that the model occasionally outputs wrong languages, particularly using Indonesian phrases or words instead of Sundanese.
 While Indonesian and Sundanese share many linguistic features, they differ significantly in morphological features (§3.2).
-This characteristic, along with the smaller amount of Sundanese data used for training LLMs, results in the models being less exposed to Sundanese words, suffixes, or prefixes compared to Indonesian.
-This leads to the occasional appearance of Indonesian words in Sundanese data, even though the sentences may be grammatically correct.
+This difference, along with the smaller amount of Sundanese data used for training LLMs, results in the models being less exposed to Sundanese words, suffixes, or prefixes compared to Indonesian.
+This leads to the occasional appearance of Indonesian words in generated Sundanese data, even though the sentences may be grammatically correct.
 5.1.3 Options and Answer Analysis
@@
 LLM_Gen 68.67% 51.00%
-Table 3: Question generation accuracy of LLM-generated datasets. The accuracy reflects the percentage of a ‘clean’ question without any major or minor errors.
+Table 4: Question generation acc. of LLM-generated datasets, measured by the % of error-free questions.
 Error Type % of questions
@@
 […]
+Wrong language 0.00% 0.00% 0.00% 6.67%
 Sent. structure 3.16% 10.13% 0.00% 0.00%
@@
 LLM_Gen 93.00% 58.67%
-Table 5: Options generation accuracy of LLM-generated datasets, measured by the % of clean/error-free options.
+Table 6: Options generation acc. of LLM-generated datasets, measured by the % of error-free options.
 Error Type % of options
@@
 […]
+Wrong language 0.00% 0.00% 0.00% 31.67%
 Sent. fluency 0.63% 0.63% 0.67% 1.33%
@@
 Options Quality
-We also evaluate the options in a similar method to the evaluation of the question.
-From Table 5, we observe that LLM_Gen generates higher quality options compared to LLM_Adapt.
+We also assess the options quality using a similar method to the question evaluation.
+From Table 6, we find LLM_Gen generates higher quality options compared to LLM_Adapt.
 However, there is still a significant gap in performance between Indonesian and Sundanese, once again highlighting the performance discrepancy between medium- and lower-resource languages.
@@
 As for LLM_Gen, the most common errors involve the presence of invalid options, particularly when no correct answers are among the options.
-For Sundanese, alongside translation errors, the model also produces some invalid options.
-However, it is important to note that the proportion of invalid options is very low for both languages, indicating that the model is still capable of generating a large number of valid options.
+For Sundanese, alongside invalid options, the model also produces some options in the wrong language.
+Still, it is important to note that the proportion of invalid options is very low for both languages, indicating that the model generates a substantial number of valid options.
 5.2 LLM vs. Human: Lexical Diversity
@@
 We also find that Human_Gen dataset has 8,596 unique tokens, higher than LLM_Gen with 6,677 unique tokens.
-Upon some sample-level analysis, we find that given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol.
+Upon sample-level analysis, we find that given the same set of categories and question concepts, humans generate more token variations that are not produced by LLMs, such as some unique terms like kalis or cimol.
 [^7] In the context of culinary, ‘kalis’ means a state of dough that is well-kneaded and ready to be processed further.
 ‘Cimol’ is an Indonesian street food made from tapioca flour.
-Although LLM can still generate some valid concepts, it tends to ask more general questions or popular concepts.
-For instance, given ‘kerupuk’ (crackers) concept, human annotators can formulate questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers).
-In contrast, LLM tends to ask more general (but still relevant) questions, like "What is the common primary ingredient of crackers?" More examples are shown in Table 12.
+Although LLM can still generate some valid data, it tends to return more popular concepts or general questions.
+For instance, given ‘kerupuk’ (crackers) concept, human annotators can create questions tied to their cultural background, such as asking about ‘kerupuk rambak’ (rambak crackers).
+In contrast, LLM tends to give more general (but still relevant) questions, like "What is the common primary ingredient of crackers?" More examples are shown in Table 12 in the Appendix.
 6 Benchmark Result
@@
 We include GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo OpenAI (2023).
-We use three prompt variations for all models (see Appendix E).
+We use three prompt variations for all models.
+[^8] Please refer to Appendix F for the prompts details.
 For a fair comparison between open and proprietary LLMs, we extract the answer key from the text generation result instead of the next token probability, using a rule-based and regex.
@@
 We use accuracy as the evaluation metric.
-Figure 3: Overall performance of LLMs on our combined test set.
-Figure 4: LLMs’ performance on LLM_Gen vs. Human_Gen dataset in Indonesian and Sundanese.
 6.2 Overall Performance
+Figure 3: LLMs’ performance on our combined test set.
 […]
-We also observe a substantial gap between Indonesian and Sundanese (10–20% accuracy drop), suggesting that current LLMs struggle to answer our Sundanese commonsense questions, even in multiple-choice settings.
-The huge gap, particularly in Merak-v4, highlights the limitations of training LLMs solely on Indonesian texts, as it does not ensure transferable performance across other local languages, especially due to the morphological differences between the languages.
-Including Sundanese texts in the training data, as demonstrated by improved performance in the Aya model, has proven effective in narrowing the performance gap between Indonesian and Sundanese.
+We also observe a substantial gap between Indonesian and Sundanese (10–20% accuracy drop), suggesting that current LLMs struggle with Sundanese questions, even in multiple-choice settings.
+This gap, particularly in Merak-v4, highlights the limitations of training LLMs solely on Indonesian texts, which does not ensure transferable performance across other local languages due to morphological differences.
+Including Sundanese texts in the training data, as shown by the improved performance in Aya, effectively narrows the performance gap between Indonesian and Sundanese.
 6.3 LLM vs. Human-Generated Data
+Figure 4: LLMs’ performance on LLM_Gen vs.
+Human_Gen in Indonesian and Sundanese.
+We combined data points from both languages for visualization, with lower quartiles typically representing Sundanese data.
 To assess LLMs’ ability to answer LLM and human-generated data, we compare their performances on our LLM_Gen and Human_Gen datasets.
 Figure 4 shows that LLMs generally perform higher on LLM_Gen, especially on MistalOrca, which is trained on English-centric data.
-This indicates that the model may possess some knowledge to answer LLM_Gen questions, even though it was not specifically trained on Indonesian or Sundanese data.
-Still, it is important to note that even in LLM_Gen, MistalOrca’s average score is still lower compared to LLMs that include Indonesian texts in their training.
-Interestingly, as the model gets stronger, we observe that the gap tends to narrow, especially in GPT models.
-For Aya, despite its relatively stronger performance compared to other open-source LLMs, its performance is significantly dropped in Human_Gen data.
-This decline is likely due to Aya being a ‘translation-heavy’ model, with a large proportion of its training data derived from translations rather than human annotations.
+This suggests the model can handle LLM_Gen questions despite not being specifically trained on Indonesian or Sundanese data.
+Still, it is important to note that MistalOrca’s average score on LLM_Gen is still lower than other LLMs that include Indonesian texts in their training.
+Interestingly, the performance gap tends to narrow as models improve, particularly for GPT models.
+For Aya, despite relatively strong performance compared to other open-source LLMs, shows significantly dropped performance in Human_Gen data.
+This decline is likely due to Aya being a ‘translation-heavy’ model Üstün et al.
+(2024), with a large proportion of its training data derived from translations rather than human annotations.
+Figure 5: LLMs performance by question category in LLM_Gen and Human_Gen for Indonesian and Sundanese.
 6.4 Performance by Question Category
@@
 Interestingly, a significant improvement is seen from GPT-4 (0613 ver.) to GPT-4 Turbo (1106 ver.), particularly in Sundanese Human_Gen.
-This suggests that the model is acquiring more knowledge, which possibly comes from the interactions on the ChatGPT web.
+This suggests that the model is “acquiring” more knowledge, possibly due to the interactions on the ChatGPT web.
 However, some categories still score below 80%, showing there is still room for improvement.
@@
 For instance, given unique activity questions such as: "What is the mandatory song to be sung during the ‘mengheningkan cipta’ (moment of silence) in the flag ceremony?" the model gives a wrong answer in open-ended generation settings even if it could identify the correct answer in the multiple-choice setup.
-More failure examples can be seen in Table 13.
+More failure examples can be seen in Table 13 in the Appendix.
 7.2 LLM-generated Data Quality
-Due to the high performance of LLMs in answering LLM-generated data, one might argue that the data is less valuable.
-Ideally, creating a larger human-generated dataset is the most straightforward method to create a high-quality dataset.
+Due to the high performance of LLMs in answering LLM-generated data, one might argue that such data is less valuable.
+Ideally, creating a larger human-generated dataset is the most straightforward method to ensure data quality.
 However, this approach may not always be feasible in the context of underrepresented languages, particularly given the limited resources.
-Building a dataset manually can be costly and challenging (we spent more than \$3,000 in total to build Human_Gen), and leveraging LLMs could be a practical and cheaper alternative, especially if we want to scale up the data.
+Building a dataset manually can be costly and challenging (we spent more than \$3,000 in total to build Human_Gen), and leveraging LLMs could be a practical and cheaper alternative, especially for scaling up the data.
 To address the limitation of LLM-generated data, we can apply additional processes to reduce the noise.
@@
 This noise can also be reduced automatically, especially in Indonesian LLM_Gen questions, where errors typically involve minor typos.
-It is also important to note that not all LLM-generated data have low quality.
-Our experiments demonstrate that LLM_Gen has significantly fewer errors compared to LLM_Adapt, particularly in Sundanese.
+It is also important to note that the quality of LLM-generated data can vary.
+Our experiments show LLM_Gen has significantly fewer errors than LLM_Adapt, particularly in Sundanese.
 We also compared the performance between the cleaned vs.
-raw version of this data (Appendix F) and found that the performance variance between the original and cleaned data of LLM_Gen is relatively minor.
+raw version of the data (Appendix G) and found that cleaning the data had a relatively minor impact on performance in LLM_Gen.
+Still, it remains crucial not to depend solely on LLMs, as they still have limitations, such as producing convincing yet incorrect outputs Pan et al.
+(2023).
 8 Conclusion
-In this study, we introduced new Indonesian and Sundanese CommonsenseQA datasets built using various methods, including automatic dataset adaptation and generation with LLMs, as well as manual dataset creation by human annotators from diverse regions.
-Our thorough analysis reveals that adaptation from the existing English data is less effective for Sundanese.
-In contrast, when directly generated in the target languages, GPT-4 Turbo can produce relevant knowledge in both languages, although the cultural depth has not yet matched the human-generated ones.
-Nevertheless, given the limited resources, combining LLMs and humans for dataset creation can be a practical solution, allowing for a more efficient dataset creation by reducing the over-reliance on a single data source.
+In this study, we introduced new Indonesian and Sundanese CommonsenseQA datasets built using various methods: automatic dataset adaptation, direct generation with LLMs, and manual dataset creation by human annotators from diverse regions.
+Our thorough analysis reveals that adapting existing English data is less effective for Sundanese.
+In contrast, direct generation in the target languages by GPT-4 Turbo produces relevant data for both languages, although the cultural depth has not yet matched the human-generated ones.
+Nevertheless, given the limited resources, combining LLMs and humans for dataset creation can be a practical solution, allowing for more efficient dataset creation by reducing the over-reliance on a single data source.
 Our goal is for LLMs to serve as beneficial tools for diverse communities, not solely those from higher-resource languages, cultures, or regions.
@@
 Nevertheless, our dataset can serve as a starting point for this line of research direction.
+Usage of GPT-4 for LLM_Gen
+LLM_Gen can initially be seen as a dataset to “only” close the gap between local LLMs and GPT-4 since we solely use GPT-4 as the data generator.
+At the time of the experiment (October–December 2023), using GPT-4 was our only sufficient option.
+Moving forward, with the rise of better LLMs that support Indonesian, such as Command-R+ [^9] https://docs.cohere.com/docs/command-r-plus or Claude 3.5 Sonnet [^10] https://www.anthropic.com/news/claude-3-5-sonnet, we can enrich the existing data by including the generation of multiple “strong” LLMs.
+This will make LLM_Gen a more comprehensive testing tool.
+Additionally, we would like to emphasize that even if LLM_Gen is seen “only” as a tool for measuring the gap between local models and GPT-4, we believe its value should not be underestimated.
+Considering the lack of strong-performing open-source LLMs in low-resource languages like Sundanese, achieving comparable performance to GPT-4 in these languages would be highly valuable for the community.
+Dataset Usage
+Human_Gen and LLM_Gen datasets should be used exclusively as test data.
+When utilizing the training data portion in the LLM_Adapt dataset, special caution is required, particularly for the Sundanese language, where numerous errors have been identified.
+Training models on this data have a risk of error propagation of incorrect information.
+Additionally, the test data portions in both LLM_Adapt and LLM_Gen should be used carefully, as they contain “silver” labels and may include some inaccuracies.
+We strongly advise using the cleaned version of the data available on our GitHub repository and accompanying the evaluation results of the LLM-generated data with those from the Human_Gen dataset.
+This is because high performance of the evaluated model, especially on the LLM-generated data, does not necessarily indicate the robustness of the model.
+One should be careful when making claims.
 Ethical Consideration
@@
 We also excluded potentially harmful questions in the LLM-generated datasets through automatic filtering.
-Our work has been reviewed by the Institutional Review Board (IRB) in our institution.
-All recruited annotators were paid above the minimum wage.
-Our datasets will be publicly available under the Creative Commons Non-Commercial (CC BY-NC 4.0) license.
+Our work has been reviewed by KAIST Institutional Review Board (IRB).
+All recruited annotators were paid above the minimum wage based on the standard in Republic of Korea.
+Our datasets are publicly available under the Creative Commons Non-Commercial (CC BY-NC 4.0) license.
+Acknowledgements
+This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.
+RS-2024-00509258, AI Guardians: Development of Robust, Controllable, and Unbiased Trustworthy AI Technology).
 Appendix A LLM-Generated Data
 A.1 Prompts Details
-A.1.1 Automatic Data Adaptation
+A.1.1 Concept Relevancy Classifier
+To classify whether a concept is relevant to the Indonesian and Sundanese context, we prompt GPT-3.5 Turbo [^11] We used gpt-3.5-turbo API checkpoint, accessed in September 2023.
+using an ensemble of five prompts.
+We condition the prompts based on whether the question concept is a verb or not, determined using the ConceptNet API.
+We set the location context as ‘Indonesia’ for Indonesian data and ‘West Java’ for Sundanese data.
+The examples of concept relevancy classifier results can be seen in Table 11.
+Below are the prompts if the question concept is a verb (e.g., skiing, studying):
+Prompt Variation 1
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Can one ’{QUESTION CONCEPT}’ in {Indonesia/West Java}? Answer with yes or no’.
+[…]
+Prompt Variation 2
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Do people in {Indonesia/West Java} familiar with ’{QUESTION CONCEPT}’ concept? Answer with yes or no’.
+[…]
+Prompt Variation 3
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Is ’{QUESTION CONCEPT}’ concept exist in {Indonesia/West Java}? Answer with yes or no’.
+[…]
+Prompt Variation 4
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+In {Indonesia/West Java}, can people ’{QUESTION CONCEPT}’? Answer with yes or no’.
+[…]
+Prompt Variation 5
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Can ’{QUESTION CONCEPT}’ be done in {Indonesia/West Java}? Answer with yes or no’.
+[…]
+Below are the prompts if the question concept is not a verb (e.g., snow, bald eagle):
+Prompt Variation 1
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Can one find ’{QUESTION CONCEPT}’ in {Indonesia/West Java}? Answer with yes or no’.
+[…]
+Prompt Variation 2
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Do people in {Indonesia/West Java} familiar with ’{QUESTION CONCEPT}’ concept? Answer with yes or no’.
+[…]
+Prompt Variation 3
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Is ’{QUESTION CONCEPT}’ concept exist in {Indonesia/West Java}? Answer with yes or no’.
+[…]
+Prompt Variation 4
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+In {Indonesia/West Java}, can people find ’{QUESTION CONCEPT}’? Answer with yes or no’.
+[…]
+Prompt Variation 5
+[…]
+[leftmargin=0pt, rightmargin=0pt]
+Text: {QUESTION}
+Concept: {QUESTION CONCEPT}
+Can ’{QUESTION CONCEPT}’ be found in {Indonesia/West Java}? Answer with yes or no’.
+[…]
+A.1.2 Automatic Data Adaptation
 There are two prompts used to adapt the CommonsenseQA into LLM_Adapt: Adapt All and Adapt Name. Below are the details of both prompts.
@@
 […]
-A.1.2 Automatic Data Generation
+A.1.3 Automatic Data Generation
 For LLM_Gen, we directly generate the data given a set of question concepts. Below are the prompt used for data generation.
@@
 Our sample-level analysis indicates that a significant number of questions generated by Merak-v4 tend to have obvious answers or are formulated as yes/no questions, such as "Apakah jagung dapat dimakan?" (eng: Is corn edible?).
-Additionally, despite being trained on Indonesian texts, the model occasionally produces questions with US-centric knowledge.
+Additionally, despite being trained on Indonesian texts, Merak-v4 occasionally produces questions with US-centric knowledge.
 For instance, it generates "Siapa yang memimpin kampanye pemilihan presiden pertama di Amerika Serikat?" (eng: Who led the first presidential election campaign in the United States?).
 A.3 Additional Analysis of the LLM-generated Data: Common Mistakes in Concept Adaptation
-Status Num (%) Concept Example
-Orig (ind) Modified (sun)
-Correct 122 (77.22%)
-hujan abu vulkanik
-(volcanic ashfall)
-hujan lebu vulkanik
-(volcanic ashfall)
-menara (tower) munara (tower)
-Wrong 36 (22.78%)
-cicak
-(house gecko)
-kadal imah
-(house lizard)
-klinik gigi
-(dental clinic)
-klinik dental
-(dental clinic)
-Table 8: Samples of Indonesian (ind) to Sundanese (sun) concept adaptation.
 Error Type Num (%) of errors
@@
 In the case of adapting concepts from ind to sun, similar to eng to ind, the majority of errors also arise from translation errors, with a larger number of errors.
-Appendix B Annotators Demographics
+Appendix B Details on Categories and Concepts
+Categories
+Our LLM_Gen and Human_Gen datasets include five selected categories, detailed as follows:
+  1. 1.
+    Culinary: Concepts in this category include everything related to culinary, starting from culinary types, cooking ingredients, cooking tools & methods, to etiquette or eating habits.
+  2. 2.
+    Places: Concepts in this category include everything related to places, starting from public facilities, landmarks, buildings, and various other concepts related to places.
+  3. 3.
+    Culture: Concepts in this category include everything related to culture, starting from cultural elements, cultural tools, cultural actors, to customs and habits that exist in Indonesia.
+  4. 4.
+    History: Concepts in this category include everything related to history, starting from historical events, historical actors, historical findings, and various other concepts related to history.
+  5. 5.
+    Activities: Concepts in this category include everything related to activities, starting from sports, hobbies, household work, and various other concepts related to activities.
+Concepts Creation
+To create the concepts, we had a human annotator manually list concepts in the selected categories, ensuring they were culturally relevant but general enough for both Indonesian and Sundanese.
+For example, in the culinary category, concepts might include ‘duck’ or ‘lime leaves,’ avoiding specific dish names.
+We reviewed the concepts and provided feedback for revisions to ensure quality.
+This process results in 150 unique concepts per category in Indonesian.
+The annotators then manually translated these concepts into Sundanese during the QA pairs creation phase (§4.2).
+We aimed for parallel concepts to ensure a fair comparison between the languages.
+Appendix C Annotators Demographics
 Data Team Ethnicity Domicile Num
@@
 The detailed annotators’ demographics are shown in Table 10.
-Appendix C Human Annotation Guideline
+Appendix D Human Annotation Guideline
 To ensure a high-quality and standardized format for question-answers annotation, we provide a specific guideline during the annotation process.
@@
     Distractor: This column only needs to be filled in when the row contains the distractor.
-Categories and Concepts
-Five selected categories are used in the data creation process.
-Each category has 150 concepts that need to be used and developed to become commonsense questions.
-The following is a list of the categories.
-  1. 1.
-    Culinary: Concepts in this category include everything related to culinary, starting from culinary types, cooking ingredients, cooking tools & methods, to etiquette or eating habits.
-  2. 2.
-    Places: Concepts in this category include everything related to places, starting from public facilities, landmarks, buildings, and various other concepts related to places.
-  3. 3.
-    Culture: Concepts in this category include everything related to culture, starting from cultural elements, cultural tools, cultural actors, to customs and habits that exist in Indonesia.
-  4. 4.
-    History: Concepts in this category include everything related to history, starting from historical events, historical actors, historical findings, and various other concepts related to history.
-  5. 5.
-    Activities: Concepts in this category include everything related to activities, starting from sports, hobbies, household work, and various other concepts related to activities.
 General Rules
-There are several rules in the data creation process. The following is the explanation.
+There are several rules in the data creation process, detailed as follows:
   1. 1.
@@
 The total number of concepts that will be accepted is 250 concepts, equivalent to the workload for each person, where the expectation is to create one question per concept.
-Appendix D Manual Data Generation Findings
+Appendix E Manual Data Generation Findings
 During the manual data generation process, we discovered several interesting findings from the data created by the annotators.
@@
 Figure 7: Answer conflict across Sundanese annotators.
-As we employ annotators from different regions, it enables us to explore variations in the data they generate.
+As we employ annotators from different regions, we can analyze variations in the data they generate.
 We examine this by calculating the number of answer conflicts that arise during the "answering question" phase of our data generation pipeline (§4.2).
-The results from Indonesian and Sundanese annotators are detailed in Figure 5 and 6, respectively.
+The results from Indonesian and Sundanese annotators are detailed in Figure 6 and 7, respectively.
 Our analysis revealed that, out of the Indonesian annotators, the one from Bali has the highest number of conflicting answers.
-However, the number of questions generated by the Bali annotator does not seem to have a lot of conflicts, suggesting that the annotator tends to generate easier questions, some of which have an obvious answer.
+However, the number of questions generated by the Bali annotator does not seem to have a lot of conflicts, suggesting that this annotator tends to generate easier questions, some of which have an obvious answer.
 This finding contrasts with the results from West Java annotators, who, despite generating questions that lead to a higher number of conflicts, do so mainly due to the creation of more challenging option distractors.
-In the case of Sundanese annotators, the variation in answer conflicts across regions is not significant.
+Interestingly, in the case of Sundanese annotators, the variation in answer conflicts across regions is not significant.
 This lack of variation can likely be attributed to the Sundanese language’s narrower geographic distribution which primarily spoken in West Java, unlike the Indonesian language, which serves as a lingua franca within the country.
-This finding points to the fact that differences in annotators’ knowledge levels may arise from various factors, one of which could be the regional culture to which they are most exposed.
-Appendix E Zero-Shot Benchmark Prompts
+However, it is important to note that given the small number of annotators, these trends cannot be conclusively linked to the annotators’ regional characteristics.
+The observed differences may also be due to individual differences rather than regional ones.
+Appendix F Zero-Shot Benchmark Prompts
 We apply three prompt variations to test LLMs’ performance on our CommonsenseQA datasets. Each prompt is described below.
@@
 Figure 8: Performance comparison of raw vs. cleaned version of LLM-generated data.
-Appendix F Effect of Synthetic Dataset Cleaning
+Appendix G Effect of Synthetic Dataset Cleaning
 To check how much cleaning the noise in synthetic (LLM-generated) data affects LLMs performance in answering our questions, we manually correct the errors in all 158 test sets of LLM_Adapt and 300 randomly sampled data from the LLM_Gen.
@@
 Table 12: Examples of data generated by GPT-4 Turbo (gpt-4-1106-preview) compared to human annotators.
-Appendix G LLM vs. Human-Generated Data Examples
+Appendix H LLM vs. Human-Generated Data Examples
 We provide some GPT-4 Turbo generation samples compared with human-generated data in Table 12.
@@
 Table 13: Examples of incorrect responses by GPT-4 Turbo (gpt-4-1106-preview) in ‘free’ or open-ended generation settings.
-Appendix H Examples from Multiple-Choice vs. ‘Free’ Generation Experiment
+Appendix I Examples from Multiple-Choice vs. ‘Free’ Generation Experiment
 We provide some GPT-4 Turbo generation samples when given the question in ‘free’ or open-ended generation settings in Table 13.
-Figure 8: LLMs performance by question category in LLM_Gen and Human_Gen for Indonesian and Sundanese.